A Complete Bibliography of Publications in Reviews of Modern Physics (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.04

Title word cross-reference

(150 < A < 190) [BR71]. 10 [PRT73b, PRT73a]. 10^{15} [GPTM78, GPTM79]. 10^{18} [GPTM78, GPTM79]. 150 < A < 190 [BR72, OWPF71]. ^{3} [Leg75, Leg76, Whe75]. A [BR71, BR72]. A > 228 [CAFE77, CAFE78]. e/h [TPL69, TPL73]. g [RW72]. γ [ABD73]. N [LG72]. p_{T} [ES77]. pp [Qui77, Qui77]. → 2π [DD79]. → 3π [DD79]. SU(2) [Act79, RS70]. SU(3) [RS70, SGM74].

-Body [LG72].

15 [Tes75]. 1976 [Sag79a, Sag79b].

A-15 [Tes75]. Abnormal [Lee75]. Above [PRT73b, PRT73a]. Absorption [VH77]. abundances [Tri75]. accelerator [HMR77]. Addenda [IIT78]. advances [She76]. algebra [Sir78a, Sir78b]. algebras [CNS75]. alignment [Ste75]. alkali [AIV77]. allowed [Hol74, Hol76]. alloys [AHM78]. American [Ano75, HCD^{+}78, LBC^{+}75, VBB^{+}75]. amorphous [AHM78]. Amplitudes [Ynd72]. Analog [AHKS72]. Analytical [Lam71]. Angular [Bri71]. Anisotropy [Vaj77]. Announcement [Ano71a, Ano72a, Ano72b, Ano74]. anomalous [CNPdR77]. any [SS79]. Application [LG72, Kad77]. Applications [BDJ^{+}72, PCF^{+}76]. applied [Hau79]. Approach [BDJ^{+}72, Hol74, NPV77, Hol76]. Aspects [LEK72]. assignments [GMRS78]. astrophysics [CW74, GP76b, Sag79a, Sag79b, GP76a]. asymptotically [SS79]. Atom [BWF73, BK71, GFK73, Ogu72b, Ogu72a, HPKY79]. Atomic [PRT73b, FTD76, PRT73a]. Atoms [Ino71, Sch73a, AIV77, IIT78]. Auger [BCF^{+}74, BCF^{+}72].

Background [Rai76, Wil79b]. Basic [EF70]. Beams [TR73a]. behavior [Fis74, Sch76, Fis75]. beta [BP78, Hol74, Hol76]. Bethe [Bro71, Ino71, IIT78]. between [PC78, ST76]. biology [Dys79]. Body [LG72, KS76]. Bond [Phi70]. Bose [CNS75, KS70]. Bosons [KS70, Qui77]. Bounds [Ynd72]. breaking [Ber74a, Ber74b, Ber75, Rey74]. breeder [Wil77]. Bremsstrahlung [Tsa77, BG70, Tsa74]. Brookhaven [HMR77, Shi74]. Brueckner [Bro71].

calculations [Day78]. Capture [ABD73, BBC^{+}77a, BBC^{+}77b, SS79]. catalysts [Sin79]. chains [LGK78]. Changing [Bet72, TR73a]. Channeling [Gem74]. Charge [BWF73, Bet72, TR73a, GMRS78, RNT75, SS79]. Charge-Changing [Bet72]. charge-independent [RNT75]. Charged [CF70, Ino71, Gem74, IIT78, Tsa74, Tsa77]. Charges [Orz70]. charm [GLR75, Ric77]. Chemical [Phi70, Tri75]. Chemistry [Deb73]. Chiral [Rey74]. Classical [Act79, Gle75, Jac77]. clusters [LS78]. Coherence [GO79, TPL69, TPL73]. Colliding [GK76]. Collision [BK71]. Collisions [Ede71, GFK73, Ino71, Ogu72a, IIT78, Qui77, Ogu72b]. Color [GMRS78]. Comment [Ano71b]. commutators [Rey74]. comparison [PC78]. Complex [VS70, VS71]. compounds [Tes75, Var76]. Compton [BG70]. Concept [Deb73]. Condensates [KS70]. Condon [Mor75]. Conduction [Kir73]. conference [HC74]. confinement [HH76]. confrontation [Jon76]. Constants [TPL69, TPL73]. constituent [ES77, Yen75]. Constraints [Ynd72]. contributions [CNPdR77]. controlled [Kap79]. Controversies [Bra70]. Convective [NPV77]. Cooperative [Hak75]. Coriolis [Ste75]. Correction [BDJ^{+}72]. corrections [Sir78a, Sir78b]. correlations [Fal79]. Cosmic [GPTM78, GP76a, GP76b, Wil79b, GPTM79]. cosmology [Fal79]. Coster [BCF^{+}74, BCF^{+}72]. Credit [Bra70]. Critical [BK71, Sag79a, Sag79b, Wil75, Fis74, Fis75, Git78, HH77, MSGH79, SGM74]. Cross [BK71, Bet72, Pow76]. Crossing [RS70]. Crystals [BL72, GRZ71, Phi70, VS70, EKL74, Gem74, McR79, SS74, Vaj77, VS71]. Current [Day78, RW72, Sir78a, Sir78b]. cycles [HCD^{+}78]. damage [Vaj77]. Data [ES77, LBGK^{+}73, RBGL^{+}71, Buc74]. dealing [PCF^{+}76]. decay [BP78, Hol74, Hol76]. decays [DD79]. decoupling [BW75]. deduced [ST76]. Defects [CB72, NP70, Mer79, Pan78]. Definitions [EF70]. Deformed [BR71, BR72]. Density [CF70, Hew75]. Dependent [LEK72]. description [Enz74, Leg75, Leg76]. Descriptions [Lam71, Gle75]. detection [HPKY79, Qui77]. Determination [TPL69, TPL73]. determinations [AIV77]. development [Sat78]. developments [Jon76]. Diatomic [Sch73b]. Dielectric [BL72]. Diffraction [Aba76, GO79]. Dilute [BG70]. dimensional [Lei75]. dimensionality [Ing75]. dipole [BF75, Fag75]. direct [Sat78]. directions [PGH^{+}74]. discovery [Jos74, Tin77]. Disordered [Dea72, BS75, EKL74]. Distribution [GP73]. Doorway [Fes74]. Double [BP78]. driven [BJ74]. dual [Sch75]. dust [Sal74]. dye [Sha75]. Dying [Sal74]. dynamic [HH77]. dynamical [GK76, LS78]. dynamics [JQ74, LK78].

early [Pai77]. Eckart [LG76]. edition [CBBG^{+}75]. Editorial [Ano71b, Ano73, Bra70, BBF^{+}70, PGH^{+}74, PCF^{+}76]. Edward [Mor75]. Effect [CF70, PRT73b, PRT73a]. effective [EO77, ST76]. Effects [BDJ^{+}72, Gem74, GO79, Hol74, Hol76, MSGH79, PC78, Sal78, Ste75, VBB^{+}75]. Eigenchannel [BBD^{+}73]. Einstein [Pai79]. Ejected [Ogu72b, Ogu72a]. Elastic [TKA72]. electric [KV79]. Electrodynamics [TPL69, TPL73]. Electroexcitation [Fag75]. electromagnetic [Tin74, Wei74]. Electron [BK71, FTD76, Gia74, GLTG71, Sch73a, Sch73b, BBC^{+}77a, BBC^{+}77b, Vaj77]. Electron-Atom [BK71]. Electron-Molecule [GLTG71]. Electron-optical [FTD76]. Electronic [Col74, McR79, AH76, Pan78]. Electrons [BG70, Mot78, Ogu72b, Ogu72a, Jac76, Pow76]. electrophysics [Sco75]. Elementary [Ede71, Mot76, Hol74, Hol76]. Elements [Bri71, Hua79, Pen79, Tri75]. embeddings [GMRS78]. Emission [GP73, DN79, VH77]. Emitted [FM73]. emphasis [Yen75]. End [Bal77, Dys79]. End-loss [Bal77]. Energies [BK71, Ter73, GPTM78, GPTM79]. Energy [Bak71, BG70, CF70, Ede71, FIM^{+}72, GP73, GLTG71, Jac70, Ogu72a, BSYP78, BSYP79, GP76a, GP76b, Igo78, VBB^{+}75, Ogu72b]. entropy [Weh78]. equation [Sch76]. equations [RNT75]. equilibrium [Fes74, Hak75]. Erratum [Ano75, BCF^{+}74, BBC^{+}77a, BSYP79, Ber74a, Ber75, BR72, CAFE78, Fis75, GPTM79, GP76a, Hol76, KRB79a, Leg76, Ogu72b, Orm75a, Par76, PRT73a, Sir78b, TPL73, Tsa77, VS71]. eV [GPTM78, GPTM79]. Evaluation [Bri71, Hua79]. evolution [KRB79a, KRB79b, LS78, Rei79]. Excitation [FM73, GLTG71, Lee75, Mot76]. Excitations [GRZ71, RNT75]. expansions [JQ74]. experiment [PC78]. Experimental [AIV77, Sco74, Whe75]. Experiments [GLTG71, GRZ71, Bah78, GO79, Jon77, Ric77]. External [VS70, VS71]. Extraterrestrial [GN71]. facilities [Wil79a]. facility [HMR77]. Fact [RGGQ74]. Factors [RW72]. fancy [RGGQ74]. Fantasies [Wil79a]. far [Hak75]. Fast [Bet72, CF70, Ino71, IIT78, Wil77]. FEED [GP73]. Fermi [CNS75, BWF73, Bro71]. Fermilab [Wil79a]. Fermion [Bak71]. Feynman [Cre78]. fiber [Sco75]. Field [Orz70, Jac77, KV79, GP73]. fields [FTD76, Mig78]. Fission [BDJ^{+}72, Spe74]. flip [Jac76]. fluid [Enz74, Ing75]. fluids [Git78, GK76, MSGH79]. Fluorescence [BCF^{+}72, BCF^{+}74]. flux [HC74]. Force [Deb73]. formalism [Hau79]. formulation [Sir78a, Sir78b]. Foundations [HvdHKN76]. Frame [HPdP78]. frames [LG76]. Free [JM73]. fuel [HCD^{+}78]. Functional [IIM75]. Functions [BL72]. Fundamental [TPL69, TPL73]. Funny [BDJ^{+}72]. Fusion [Rib75, BJ74]. future [Wil79a].

galaxies [Jon76]. Galaxy [Fal79]. gas [MSGH79]. gas-liquid [MSGH79]. Gaseous [Bet72]. Gases [BG70]. gauge [Ber74a, Ber74b, Ber75, Cre78, GMRS78, Kog79, Sir78a, Sir78b, Wei74]. General [Weh78, HvdHKN76]. Generators [Orz70]. giant [BF75]. glass [Mot78, Ols79]. globular [LS78]. Graded [CNS75]. gradient [KV79]. Graph [EF70]. Graphical [Bri71, Hua79]. Gravity [MSGH79, GO79]. Ground [Bak71]. Ground-State [Bak71]. Group [LBGK^{+}73, LG72, kM73, Ano75, Fis74, Fis75, Kad77, Wil75, RBGL^{+}71]. guide [EO77].

hadron [KS76]. hadronic [BSYP78, BSYP79, DD79, Yen75]. Hadrons [SGM74, Aba76]. Heavy [Bet72, Gle75, GK76, Qui77]. heavy-ion [Gle75]. Helium [GRZ71, TKA72, Col74, Mar77]. Higgs [Ber74a, Ber74b, Ber75]. High [BG70, Coh77, Ede71, FIM^{+}72, Jac70, Sor73, Ter73, BSYP78, BSYP79, GP76a, GP76b, Hew75, Igo78, SS79]. High-Energy [BG70, FIM^{+}72, Jac70, BSYP78, BSYP79, GP76a, GP76b, Igo78]. High-level [Coh77]. Hills [BDJ^{+}72]. homogeneous [KRB79a, KRB79b]. hydrodynamic [Enz74]. Hydrodynamics [Git78]. Hydrogen [TR73a]. hyperfine [AIV77].

II [JM73, Rei79, Sag79b]. Impact [FM73, Sch73a, Sch73b, SS79]. Implications [TPL69, TPL73]. impurities [Pan78]. independent [RNT75]. induced [Hau79]. Inelastic [Ino71, IIT78]. Inertia [Sor73]. inhomogeneous [Rei79]. Inner [GFK73, Pow76]. Inner-Shell [GFK73, Pow76]. instability [NPV77, Tes75]. Insulators [CB72]. Integral [Bri71]. Integrated [Tie77]. Interacting [GMM71]. Interaction [KRB79a, LGK78, ST76, KRB79b, Rei79]. interactions [BSYP78, Gle75, KRB79a, KRB79b, Mar77, Rei79, Sir78a, Wei74, BSYP79, Sir78b]. interference [Sal78]. intermediate [Igo78, Qui77]. intermediate-and [Igo78]. Interplay [Gle75]. Interpretation [Bal70, Hau79, KS76]. intersecting [HMR77]. Interstellar [Wat76]. Introduction [Lei75, kM73, Kog79, Sch75]. introductory [EO77]. invariant [RNT75]. inverse [Hau79]. Inversion [Buc74]. investigated [LK78]. invitation [Pin77]. Ion [GFK73, Ogu72b, Ogu72a, Gle75]. Ion-Atom [GFK73, Ogu72b, Ogu72a]. Ionicity [Phi70]. Ionization [CF70, HPKY79, Pow76]. Ions [Bet72, GK76]. ISABELLE [HMR77]. Isobaric [AHKS72, BK79].

J [Tin77]. journey [Esa74].

keV [PRT73b, PRT73a]. key [Van78]. Kinds [KS70]. Kondo [Wil75]. Kronig [BCF^{+}74, BCF^{+}72].

laboratory [PC78, HMR77]. Landau [Bro71]. Large [ES77, Ryl75]. Laser [BJ74]. Laser-driven [BJ74]. lasers [Sha75]. lattice [Cre78, Kog79]. layers [Sag79b]. Lectures [Sag79a, Sag79b]. Lepton [RW72]. leptons [Qui77, Tsa74, Tsa77]. level [Coh77]. Levels [OWPF71]. Lie [CNS75]. Light [FM73, Ano75, Coh77, LK78, LBC^{+}75, SS79]. light-water [Ano75, Coh77, LBC^{+}75]. Limits [GN71]. linebreadths [VH77]. lines [DN79]. links [Lik79]. liquid [BH76, Col74, Leg75, Leg76, Mar77, MSGH79, SS74, Wil77]. liquids [LK78]. loading [DG77]. Local [And78]. localized [And78]. Long [Esa74]. Loss [CF70, Bal77]. Low [BK71, GLTG71].

machines [Bal77]. Macroscopic [TPL69, TPL73, Sch76]. Magnetic [CB72, HC74, TKA72, CNPdR77, Fag75]. magnetism [Van78]. management [HCD^{+}78]. Many [Bak71]. Many-Fermion [Bak71]. Mass [GN71, CAFE77, CAFE78]. master [Sch76]. materials [VBB^{+}75]. mathematics [CNS75]. Matrices [RS70]. Matrix [Bri71, Hua79]. Matter [CF70, BH76, Day78, Lie76, Mig78, PW79]. meaning [Jac77]. Measurements [BF75, Orm75b, Orm75a]. Mechanics [Bal70, Van78]. mechanism [AHM78, Ber74a, Ber74b, Ber75]. Mechanisms [SS79]. Media [Bet72, Mer79]. Medium [Lam71, KRB79b, Rei79, KRB79a]. meson [Rey74]. meson-nucleon [Rey74]. metal [Sin79, Vaj77, Wil77]. metals [KV79]. Method [BBD^{+}73]. Methods [LG72, IIM75]. microscopic [Sch76]. microwave [Wil79b]. Migdal [Bro71]. Mills [Act79]. mirror [Bal77]. Mixed [Var76]. Mixed-valence [Var76]. Mod [Ano75]. mode [Sco74]. model [Rai76]. Models [GMM71, Jac70, SY71, ES77, Sch75]. Modes [BL72, Mot76]. molecular [Buc74]. Molecule [GLTG71, Wat76]. Molecules [Ino71, Sch73b, HPdP78, IIT78, LG76]. Moment [Sor73, CNPdR77]. moments [And78]. monoenergetic [BF75]. motion [Boh76, Gem74, Ing75, RNT75]. Multiparticle [FIM^{+}72]. Multiple [JQ74]. multipole [HPdP78]. muon [CNPdR77].

Narrow [SY71]. National [HMR77]. Natural [Dav72]. near [Git78, MSGH79]. nerve [Sco75]. Network [Sch76]. Neutrino [TR73b, Bah78, RGGQ74]. Neutron [ABD73, Shi74, GO79]. Neutron-Capture [ABD73]. NMR [GRZ71]. No [Ano75]. noncubic [KV79]. Nonlinear [GMM71, PC78, KRB79a, KRB79b, Rei79, Sag79a, She76]. nonphysical [Hak75]. Nonrotational [BR71, BR72]. Nonspherical [OWPF71]. Nuclear [ABD73, BBD^{+}73, BDJ^{+}72, JM73, Sor73, Spe74, Day78, Fag75, HCD^{+}78, Lee75, Mig78, PW79, Rai76, Sat78, ST76]. Nuclei [BR71, BR72, GK76, OWPF71, BK79, Boh76, EO77, Ste75]. nucleon [Rey74]. nucleons [ST76]. nucleus [BBC^{+}77b, Igo78, Mot76, BBC^{+}77a]. Numerical [Dea72].

objects [DN79]. observation [Jon76]. Odd [BR71, BR72]. Odd- [BR71, BR72]. one [HPKY79]. one-atom [HPKY79]. open [Dys79]. operators [EO77]. Oppenheimer [Sag79a, Sag79b]. Optical [BS75, Hap72, Lam71, FTD76, Ols79, Tie77]. optics [She76, Tie77]. Orbital [BBC^{+}77b, BBC^{+}77a]. Orbitals [Dav72]. ordered [Enz74, Mer79]. origin [GP76a, GP76b, Jon76, Pen79, Tri75]. Orthogonal [LG72]. Oscillation [BWF73]. Other [GMM71, Pan78]. our [Pin77]. Outline [Sat78]. outlook [Aba76].

Pair [Tsa74, Tsa77]. Particle [LBGK^{+}73, OWPF71, RBGL^{+}71, Ter73, CAFE77, CAFE78, GPTM78, GPTM79, Hol74, Hol76, JQ74, Par76, SS79, Tin77, RBGL^{+}71]. Particles [CF70, Ede71, Ino71, Gem74, IIT78]. Penetrating [Bet72]. Percolation [Kir73]. personal [Tin77]. perspective [VH77]. Perturbation [Bak71, LEK72, IIM75]. Phase [DG77, TPL69, TPL73, Sco74, Shi74]. phases [Leg75, Leg76]. phenomena [Hak75, HH77, Tie77, Wil75]. Phonon [Mar77]. Phonon-phonon [Mar77]. Photoabsorption [BWF73]. Photoelectric [PRT73b, PRT73a]. Photon [GN71, Ter73, BSYP78, BSYP79, Sal78, Yen75]. photons [BF75]. Phys [Ano75]. Physical [Ano75, Hau79, HCD^{+}78, LBC^{+}75, TPL69, TPL73, VBB^{+}75, EKL74]. physicist [NPV77]. Physics [Dys79, Sha75, SS74, Wil77, CNS75, ES77, GPTM78, GPTM79, Hew75, LGK78, PCF^{+}76, RGGQ74, VBB^{+}75]. picosecond [LK78]. Pion [Mig78, Yen75]. Plasma [Kap79, HH76, Sag79a, Sag79b]. plasmas [PC78]. point [Git78, MSGH79, Pan78]. Polariton [BL72]. Polarization [FM73, Jac76]. Policy [BBF^{+}70]. polyatomic [HPdP78, LG76]. polymer [LGK78]. Pomeron [BW75]. Populations [GMM71]. Potentials [FLS71]. power [Ryl75]. pre [Fes74]. pre-equilibrium [Fes74]. Probabilities [BCF^{+}72, BCF^{+}74]. Problem [LG72, TR73b, Wil75]. problems [GP76a, GP76b, LGK78, Sag79a, Sag79b, VBB^{+}75]. Process [BDJ^{+}72]. Processes [Jac70, TR73a, Ter73, Bal77]. Production [GFK73, Qui77, Ter73, Tsa74, Tsa77]. Progress [TR73b, Wei74]. Propagation [Lam71, Ols79]. Properties [BGDP^{+}70, CB72, Dav72, Dea72, LBGK^{+}73, RBGL^{+}71, TKA72, BS75, BSYP78, BSYP79, CBBG^{+}75, EKL74, FTD76, Tin74, TBGK^{+}76, Weh78, Whe75, Par76]. proposal [Rai76]. prospects [HvdHKN76]. Proton [HMR77, GMRS78, Igo78]. proton-nucleus [Igo78]. Proton-proton [HMR77]. psi [Ric77]. Pulsars [Hew75]. Pulse [Lam71]. pulses [LK78]. Pumping [Hap72]. puzzles [Pai77].

quantal [Gle75]. Quantum [Bal70, GRZ71, Jac77, Orz70, Sal78, TPL69, TPL73, Van78, Pai79]. quark [Jon77]. quarks [Kad77]. quartets [BK79]. quasars [DN79]. Quasi [JM73]. Quasi-Free [JM73].

Radiation [BG70, VBB^{+}75, Jac76, Vaj77, Wil79b]. radiative [Sir78a, Sir78b]. Radio [Ryl75]. radioactive [Coh77]. Radioactivity [Pai77]. Raman [BL72]. randomly [EKL74]. Ray [BCF^{+}72, GPTM78, GPTM79, BCF^{+}74]. Rays [ABD73, GP76a, GP76b]. reaction [Kap79]. Reactions [BBD^{+}73, FIM^{+}72, Fes74, KS76, Sat78, Wat76]. reactor [Ano75, LBC^{+}75, Rib75, Wil77]. reactors [Coh77]. readers [Pin77]. reborn [Sal74]. Recoil [Hol74, Hol76]. recollection [Tin77]. reconnection [Sag79b]. Region [OWPF71, CAFE77, CAFE78]. regularization [Lei75]. Related [BBD^{+}73, EKL74, Gem74]. relating [VBB^{+}75]. Relations [Ynd72, HPdP78]. relativistic [Hua79]. Relativity [HvdHKN76]. relaxation [Orm75a, Orm75b]. Renormalization [kM73, Fis74, Fis75, Kad77, Wil75]. Report [HCD^{+}78, LBC^{+}75, TR73b, VBB^{+}75, Ano75]. Representation [Bri71]. research [Igo78]. resolving [Ryl75]. Resonance [HPKY79, SY71, BF75]. Resonances [AHKS72, Sch73a, Sch73b, McR79]. Resonant [Lam71]. Response [BL72]. Responsibility [Bra70]. Rev [Ano75]. Review [BGDP^{+}70, BK71, CBBG^{+}75, LBGK^{+}73, RBGL^{+}71, SY71, TBGK^{+}76, CNPdR77, Jon76, Jon77, Rey74, SGM74, Par76]. Reviewer [Bra70]. Reviews [PCF^{+}76]. Revisited [Ino71, IIT78]. Rigorous [Ynd72]. rings [Jac76]. rotating [Ing75]. rotation [Ste75]. Rotational [Boh76, GLTG71]. rotationally [RNT75]. rules [Cre78].

safety [Ano75, LBC^{+}75, Wil77]. Scattering [BL72, BG70, GLTG71, JM73, Ynd72, Aba76, Buc74, Hau79, JQ74, Shi74]. Search [GLR75, Jon77]. Sections [BK71, Bet72, Pow76]. self [Hau79]. self-induced [Hau79]. semiconductors [Pan78]. semihistorical [VH77]. Series [Bak71]. several [JQ74]. Shell [BDJ^{+}72, GFK73, Pow76]. Shell-Correction [BDJ^{+}72]. shock [DG77]. shock-wave [DG77]. showers [GPTM78, GPTM79]. sigma [Rey74]. similar [DN79]. Single [OWPF71, CAFE77, CAFE78]. Single-Particle [OWPF71, CAFE77]. Singular [FLS71, Sag79b]. Singularity [Bak71]. slow [Ing75]. Society [Ano75, HCD^{+}78, LBC^{+}75, VBB^{+}75]. Soft [Sco74]. Soft-mode [Sco74].

Solar [Bah78, TR73b]. Solid [Bet72, GRZ71, TKA72, AH76]. Solids [NP70, BS75, LK78]. solutions [Act79]. Some [EF70, GP76a, GP76b, Igo78, LGK78, PGH^{+}74]. Space [KRB79b, Rei79, KRB79a]. Spectra [Ogu72b, Ogu72a, ST76]. spectroscopy [HPKY79, Sal78, Sco74]. spheroidal [Rai76]. Spin [Sor73, HvdHKN76, Jac76, Kog79]. spin-flip [Jac76]. Spontaneous [Ber74b, Ber74a, Ber75]. stability [GMRS78, Lie76]. stars [Sal74]. State [Bak71, Day78]. States [Bet72, BR71, BR72, NP70, And78, BH76, CAFE77, CAFE78, Col74, Fes74, Lee75]. Statistical [Bal70, LGK78]. Status [RW72]. storage [HMR77, Jac76]. strings [Kad77, Sch75]. strong [Wei74]. Structural [Tes75, Sco74, Shi74]. Structure [Bak71, JM73, Sin79, AH76, AIV77, Pan78, Yen75]. structures [HC74]. Studies [ABD73, Dea72, BS75, Sco74, Shi74]. study [Ano75]. summary [HC74]. Superconducting [Lik79]. superconductivity [Gia74, Tes75]. Superconductors [TPL69, TPL73, HC74, Tin74]. supercurrents [Jos74]. superfluid [Whe75]. Suppl [Ano75]. Supplement [CBBG^{+}75]. surface [Col74, McR79]. surfaces [AH76]. Survey [BR71, BR72, CAFE77, GRZ71, CAFE78]. switching [AHM78]. symmetric [HPdP78]. Symmetry [Orz70, Ber74a, Ber74b, Ber75, CNS75, Rey74]. Synchrotron [BG70, Jac76]. System [Bak71]. Systems [Bro71, Dea72, EKL74, Enz74, Hak75, HH76, Kog79, Rib75, RNT75, Sch76].

technique [Lei75]. techniques [Kad77]. technologies [VBB^{+}75]. telescopes [Ryl75]. Tensor [RNT75]. Terrestrial [GN71]. their [Jon76]. theme [PW79]. Theorems [Ede71, BW75]. theoretical [Aba76, CNPdR77, Jon76, Leg75, Leg76]. Theories [BBD^{+}73, Bro71, Act79, Ber74a, Ber74b, Ber75, GMRS78, Orm75a, Orm75b, Sir78a, Sir78b, Wei74]. Theory [AHKS72, BL72, EF70, GLTG71, HH76, HH77, Ino71, LEK72, Orz70, Cre78, EKL74, Fis74, Fis75, IIM75, IIT78, Jac77, Kog79, Mer79, Pai79, PC78, Sat78, Sch75, Sch76]. thermal [Hak75]. Thermodynamic [TKA72]. Thermonuclear [CW74, Kap79]. Thomas [BWF73]. three [KRB79a, KRB79b, Rei79]. three-wave [KRB79a, KRB79b, Rei79]. threshold [AHM78]. Time [Dys79, LEK72, KRB79a, KRB79b, Rei79]. Time-Dependent [LEK72]. topological [Mer79]. toroidal [HH76]. torsion [HvdHKN76]. Total [BK71]. transfer [SS79]. transformation [HPdP78]. Transformations [Orz70]. Transition [BCF^{+}72, BCF^{+}74]. transitions [DG77, Fag75, HPdP78, Sco74, Shi74]. transparency [Hau79]. transport [HH76]. transverse [Jac76]. Traversing [BG70]. Tunneling [NP70, Esa74, Gia74]. tunnelling [Jos74]. Turbulence [Sag79a]. Two [Enz74, KS70, Ter73, Ing75, KS76, Pai77, Sal78, Yen75]. two-body [KS76]. two-dimensionality [Ing75]. Two-fluid [Enz74]. Two-Photon [Ter73, Sal78]. two-pion [Yen75].

Uhler [Mor75]. Ultrashort [Lam71]. understanding [Jac76, Van78, BH76]. unified [GMRS78]. Unitary [LG72]. universality [Sir78a, Sir78b]. universe [Dys79]. Uses [Dav72]. Using [TPL69, TPL73].

Vacancy [GFK73]. vacuum [Lee75]. valence [Var76]. Variations [PW79]. vector [Qui77]. vectors [LG76]. velocities [SS79]. Vibrational [BL72, Dea72, LK78, Orm75b, BS75, Orm75a]. Vibrations [VS70, VS71]. Volterra [GMM71]. volume [LGK78].

waste [Coh77, HCD^{+}78]. water [Ano75, Coh77, LBC^{+}75]. wave [DG77, KRB79a, KRB79b, PC78, Rei79, Tie77]. waveguides [Ols79, Tie77]. waves [Sag79a]. Weak [DD79, Lik79, Sir78a, Sir78b, Wei74]. without [Dys79]. Work [SY71].

X [BCF^{+}72, BCF^{+}74]. X-ray [BCF^{+}74, BCF^{+}72].

Yang [Act79]. Yields [BCF^{+}72, BCF^{+}74].

References

Abarbanel:1976:DSH

[Aba76] Henry D. I. Abarbanel. Diffraction scattering of hadrons: The theoretical outlook. Reviews of Modern Physics, 48 (3):435–465, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

            link.aps.org/doi/10.1103/RevModPhys.48.435;   http://
            rmp.aps.org/abstract/RMP/v48/i3/p435_1.

Arad:1973:NSN

[ABD73] B. Arad and G. Ben-David. Nuclear studies with neutron-capture γ rays. Reviews of Modern Physics, 45(2):230–272, April 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

            tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
            RevModPhys.45.230;   http://rmp.aps.org/abstract/RMP/
            v45/i2/p230_1.

Actor:1979:CSY

[Act79] Alfred Actor. Classical solutions of SU(2) Yang–Mills theories. Reviews of Modern Physics, 51(3):461–525, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

            tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
            RevModPhys.51.461;   http://rmp.aps.org/abstract/RMP/
            v51/i3/p461_1.

Appelbaum:1976:ESS

[AH76] Joel A. Appelbaum and D. R. Hamann. The electronic structure of solid surfaces. Reviews of Modern Physics, 48 (3):479–496, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.479;   http://
rmp.aps.org/abstract/RMP/v48/i3/p479_1.

Auerbach:1972:TIA

[AHKS72] Naftali Auerbach, Jörg Hüfner, A. K. Kerman, and C. M. Shakin. A theory of isobaric analog resonances. Reviews of Modern Physics, 44(1):48–125, January 1972. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
44.48;   http://rmp.aps.org/abstract/RMP/v44/i1/p48_1.

Adler:1978:MTS

[AHM78] David Adler, Heinz K. Henisch, and Sir Nevill Mott. The mechanism of threshold switching in amorphous alloys. Reviews of Modern Physics, 50(2):209–220, April 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.209;   http:
//rmp.aps.org/abstract/RMP/v50/i2/p209_1.

Arimondo:1977:EDH

[AIV77] E. Arimondo, M. Inguscio, and P. Violino. Experimental determinations of the hyperfine structure in the alkali atoms. Reviews of Modern Physics, 49(1):31–75, January 1977. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
49.31;   http://rmp.aps.org/abstract/RMP/v49/i1/p31_1.

Anderson:1978:LML

[And78] P. W. Anderson. Local moments and localized states. Reviews of Modern Physics, 50(2):191–201, April 1978. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.191;   http://rmp.aps.org/abstract/RMP/
v50/i2/p191_1.

Anonymous:1971:A

[Ano71a] Anonymous. Announcement. Reviews of Modern Physics, 43(1):98, January 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.43.98;   http://
rmp.aps.org/abstract/RMP/v43/i1/p98_1.

Anonymous:1971:EC

[Ano71b] Anonymous. Editorial comment. Reviews of Modern Physics, 43(4):641, October 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.43.641;   http://
rmp.aps.org/abstract/RMP/v43/i4/p641_1.

Anonymous:1972:Aa

[Ano72a] Anonymous. Announcement. Reviews of Modern Physics, 44(1):126, January 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.126;   http://
rmp.aps.org/abstract/RMP/v44/i1/p126_2.

Anonymous:1972:Ab

[Ano72b] Anonymous. Announcement. Reviews of Modern Physics, 44(2):450, April 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.450;   http://
rmp.aps.org/abstract/RMP/v44/i2/p450_1.

Anonymous:1973:E

[Ano73] Anonymous. Editorial. Reviews of Modern Physics, 45 (2):326, April 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.45.326;   http://
rmp.aps.org/abstract/RMP/v45/i2/p326_1.

Anonymous:1974:A

[Ano74] Anonymous. Announcement. Reviews of Modern Physics, 46(2):430, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.430;   http://
rmp.aps.org/abstract/RMP/v46/i2/p430_1.

Anonymous:1975:ERA

[Ano75] Anonymous. Erratum: Report to the American Physical Society by the study group on light-water reactor safety [Rev. Mod. Phys. 47, Suppl. No. 1 (1975)]. Reviews of Modern Physics, 47(4):979, October 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.979;   http://
rmp.aps.org/abstract/RMP/v47/i4/p979_1. See [LBC^{+}75].

Bahcall:1978:SNE

[Bah78] John N. Bahcall. Solar neutrino experiments. Reviews of Modern Physics, 50(4):881–903, October 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.881;   http:
//rmp.aps.org/abstract/RMP/v50/i4/p881_1.

Baker:1971:SSP

[Bak71] George A. Baker, Jr. Singularity structure of the perturbation series for the ground-state energy of a many-fermion system. Reviews of Modern Physics, 43(4):479–531, October 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.43.479;   http://rmp.aps.org/abstract/RMP/
v43/i4/p479_1.

Ballentine:1970:SIQ

[Bal70] L. E. Ballentine. The statistical interpretation of quantum mechanics. Reviews of Modern Physics, 42(4):358– 381, October 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.42.358;   http://
rmp.aps.org/abstract/RMP/v42/i4/p358_1.

Baldwin:1977:ELP

[Bal77] D. E. Baldwin. End-loss processes from mirror machines. Reviews of Modern Physics, 49(2):317–339, April 1977. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.317;   http://rmp.aps.org/abstract/RMP/
v49/i2/p317_1.

Bambynek:1977:EOE

[BBC^{+}77a] W. Bambynek, H. Behrens, M. H. Chen, B. Crasemann, M. L. Fitzpatrick, K. W. D. Ledingham, H. Genz, M. Mutterer, and R. L. Intemann. Erratum: Orbital electron capture by the nucleus. Reviews of Modern Physics, 49(4):961–962, October 1977.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.961;   http://rmp.aps.org/abstract/RMP/

v49/i4/p961_1. See [BBC^{+}77b].

Bambynek:1977:OEC

[BBC^{+}77b] W. Bambynek, H. Behrens, M. H. Chen, B. Crasemann, M. L. Fitzpatrick, K. W. D. Ledingham, H. Genz, M. Mutterer, and R. L. Intemann. Orbital electron capture by the nucleus. Reviews of Modern Physics, 49(1):77–221, January 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
49.77;   http://rmp.aps.org/abstract/RMP/v49/i1/p77_1.

See erratum [BBC^{+}77a].

Barrett:1973:EMR

[BBD^{+}73] R. F. Barrett, L. C. Biedenharn, Michael Danos, P. P. Delsanto, W. Greiner, and H. G. Wahsweiler. The eigenchannel method and related theories for nuclear reactions. Reviews of Modern Physics, 45(1):44–108, January 1973. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
45.44;   http://rmp.aps.org/abstract/RMP/v45/i1/p44_1.

Branscomb:1970:EP

[BBF^{+}70] L. M. Branscomb, S. J. Buchsbaum, U. Fano, J. D. Jackson, A. K. Kerman, and J. Krumhansl. Editorial policy. Reviews of Modern Physics, 42(4):357, October 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.42.357;   http:
//rmp.aps.org/abstract/RMP/v42/i4/p357_1.

Bambynek:1972:XRF

[BCF^{+}72] Walter Bambynek, Bernd Crasemann, R. W. Fink, H.-U. Freund, Hans Mark, C. D. Swift, R. E. Price, and P. Venugopala Rao. X-ray fluorescence yields, Auger, and Coster– Kronig transition probabilities. Reviews of Modern Physics, 44 (4):716–813, October 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.44.716;   http:
//rmp.aps.org/abstract/RMP/v44/i4/p716_1. See erratum

[BCF^{+}74].

Bambynek:1974:EXR

[BCF^{+}74] Walter Bambynek, Bernd Crasemann, R. W. Fink, H. U. Freund, Hans Mark, C. D. Swift, R. E. Price, and P. Venugopala Rao. Erratum: X-ray fluorescence yields, Auger, and Coster– Kronig transition probabilities. Reviews of Modern Physics, 46(4):853, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.853;   http://
rmp.aps.org/abstract/RMP/v46/i4/p853_1. See [BCF^{+}72].

Brack:1972:FHS

[BDJ^{+}72] M. Brack, Jens Damgaard, A. S. Jensen, H. C. Pauli, V. M. Strutinsky, and C. Y. Wong. Funny hills: The shell-correction approach to nuclear shell effects and its applications to the fission process. Reviews of Modern Physics, 44(2):320–405, April 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.320;   http://rmp.aps.org/abstract/RMP/
v44/i2/p320_1.

Bernstein:1974:ESS

[Ber74a] Jeremy Bernstein. Erratum: Spontaneous symmetry breaking, gauge theories, the Higgs mechanism and all that. Reviews of Modern Physics, 46(4):855, October 1974. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.855;   http://rmp.aps.org/abstract/RMP/

v46/i4/p855_1. See [Ber74b, Ber75].

Bernstein:1974:SSB

[Ber74b] Jeremy Bernstein. Spontaneous symmetry breaking, gauge theories, the Higgs mechanism and all that. Reviews of Modern Physics, 46(1):7–48, January 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.7;   http://
rmp.aps.org/abstract/RMP/v46/i1/p7_1. See errata [Ber74a,

Ber75].

Bernstein:1975:ESS

[Ber75] Jeremy Bernstein. Erratum: Spontaneous symmetry breaking, gauge theories, the Higgs mechanism and all that. Reviews of Modern Physics, 47(1):259, January 1975. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.259;   http://rmp.aps.org/abstract/RMP/

v47/i1/p259_1. See [Ber74b, Ber74a].

Betz:1972:CSC

[Bet72] Hans-Dieter Betz. Charge states and charge-changing cross sections of fast heavy ions penetrating through gaseous and solid media. Reviews of Modern Physics, 44(3):465–539, July 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.465;   http://rmp.aps.org/abstract/RMP/
v44/i3/p465_1.

Berman:1975:MGD

[BF75] B. L. Berman and S. C. Fultz. Measurements of the giant dipole resonance with monoenergetic photons. Reviews of Modern Physics, 47(3):713–761, July 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.47.713;   http:
//rmp.aps.org/abstract/RMP/v47/i3/p713_1.

Blumenthal:1970:BSR

[BG70] George R. Blumenthal and Robert J. Gould. Bremsstrahlung, synchrotron radiation, and Compton scattering of high-energy electrons traversing dilute gases. Reviews of Modern Physics, 42(2):237–270, April 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.42.237;   http://
rmp.aps.org/abstract/RMP/v42/i2/p237_1.

Barbaro-Galtieri:1970:RPP

[BGDP^{+}70] Angela Barbaro-Galtieri, Stephen E. Derenzo, Leroy R. Price, Alan Rittenberg, Arthur H. Rosenfeld, Naomi Barash-Schmidt, Claude Bricman, Matts Roos, Paul Söding, and Charles G. Wohl (Particle Data Group). Review of particle properties. Reviews of Modern Physics, 42(1):87–200, January 1970. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
42.87;   http://rmp.aps.org/abstract/RMP/v42/i1/p87_1.

Barker:1976:WLU

[BH76] J. A. Barker and D. Henderson. What is “liquid”? Understanding the states of matter. Reviews of Modern Physics, 48 (4):587–671, October 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.587;   http://
rmp.aps.org/abstract/RMP/v48/i4/p587_1.

Brueckner:1974:LDF

[BJ74] Keith A. Brueckner and Siebe Jorna. Laser-driven fusion. Reviews of Modern Physics, 46(2):325–367, April 1974. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.325;   http://rmp.aps.org/abstract/RMP/
v46/i2/p325_1.

Bederson:1971:TEA

[BK71] Benjamin Bederson and L. J. Kieffer. Total electron-atom collision cross sections at low energies — a critical review. Reviews of Modern Physics, 43(4):601–640, October 1971. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.43.601;   http://rmp.aps.org/abstract/RMP/
v43/i4/p601_1.

Benenson:1979:IQN

[BK79] Walter Benenson and Edwin Kashy. Isobaric quartets in nuclei. Reviews of Modern Physics, 51(3):527–540, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.527;   http://rmp.aps.org/abstract/RMP/
v51/i3/p527_1.

Barker:1972:RFT

[BL72] A. S. Barker, Jr. and R. Loudon. Response functions in the theory of Raman scattering by vibrational and polariton modes in dielectric crystals. Reviews of Modern Physics, 44(1):18–47, January 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.44.18;   http://rmp.aps.org/abstract/
RMP/v44/i1/p18_1.

Bohr:1976:RMN

[Boh76] Aage Bohr. Rotational motion in nuclei. Reviews of Modern Physics, 48(3):365–374, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.48.365;   http:
//rmp.aps.org/abstract/RMP/v48/i3/p365_1.

Bryman:1978:DBD

[BP78] Douglas Bryman and Charles Picciotto. Double beta decay. Reviews of Modern Physics, 50(1):11–21, January 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
50.11;   http://rmp.aps.org/abstract/RMP/v50/i1/p11_1.

Bunker:1971:SNS

[BR71] M. E. Bunker and C. W. Reich. A survey of nonrotational states of deformed odd-A nuclei (150 < A < 190). Reviews of Modern Physics, 43(3):348–423, July 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.43.348;   http:
//rmp.aps.org/abstract/RMP/v43/i3/p348_1. See erratum

[BR72].

Bunker:1972:ESN

[BR72] M. E. Bunker and C. W. Reich. Erratum: A Survey of Nonrotational States of Deformed Odd-A Nuclei 150 < A < 190. Reviews of Modern Physics, 44(1):126, January 1972. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.126;   http://rmp.aps.org/abstract/RMP/
v44/i1/p126_1. See [BR71].

Branscomb:1970:ECC

[Bra70] Lewis M. Branscomb. Editorial: Credit controversies and the Reviewer’s responsibility. Reviews of Modern Physics, 42 (3):289, July 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.42.289;   http://
rmp.aps.org/abstract/RMP/v42/i3/p289_1.

Briggs:1971:EME

[Bri71] John S. Briggs. Evaluation of matrix elements from a graphical representation of the angular integral. Reviews of Modern Physics, 43(2):189–230, April 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.43.189;   http:
//rmp.aps.org/abstract/RMP/v43/i2/p189_1.

Brown:1971:LBB

[Bro71] G. E. Brown. Landau, Brueckner–Bethe, and Migdal theories of Fermi systems. Reviews of Modern Physics, 43 (1):1–14, January 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.43.1;   http://
rmp.aps.org/abstract/RMP/v43/i1/p1_1.

Barker:1975:OSV

[BS75] A. S. Barker, Jr. and A. J. Sievers. Optical studies of the vibrational properties of disordered solids. Reviews of Modern Physics, 47(S2):S1–S179, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.47.S1;   http://
rmp.aps.org/abstract/RMP/v47/iS2/pS1_2.

Bauer:1978:HPP

[BSYP78] T. H. Bauer, R. D. Spital, D. R. Yennie, and F. M. Pipkin. The hadronic properties of the photon in high-energy interactions. Reviews of Modern Physics, 50(2):261–436, April 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.261;   http://rmp.aps.org/abstract/RMP/

v50/i2/p261_1. See erratum [BSYP79].

Bauer:1979:EHP

[BSYP79] T. H. Bauer, R. D. Spital, D. R. Yennie, and F. M. Pipkin. Erratum: The hadronic properties of the photon in high-energy interactions. Reviews of Modern Physics, 51(2):407–409, April 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.407;   http://rmp.aps.org/abstract/RMP/
v51/i2/p407_1. See [BSYP78].

Buck:1974:IMS

[Buc74] U. Buck. Inversion of molecular scattering data. Reviews of Modern Physics, 46(2):369–389, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.369;   http:
//rmp.aps.org/abstract/RMP/v46/i2/p369_1.

Brower:1975:PDT

[BW75] Richard C. Brower and J. H. Weis. Pomeron decoupling theorems. Reviews of Modern Physics, 47(3):605–635, July 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.605;   http://rmp.aps.org/abstract/RMP/
v47/i3/p605_1.

Ball:1973:PCO

[BWF73] J. A. Ball, J. A. Wheeler, and E. L. Firemen. Photoabsorption and charge oscillation of the Thomas–Fermi atom. Reviews of Modern Physics, 45(3):333–352, July 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.45.333;   http:
//rmp.aps.org/abstract/RMP/v45/i3/p333_1.

Chasman:1977:SSP

[CAFE77] R. R. Chasman, I. Ahmad, A. M. Friedman, and J. R. Erskine. Survey of single-particle states in the mass region A > 228. Reviews of Modern Physics, 49(4):833–891, October 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.833;   http://rmp.aps.org/abstract/RMP/

v49/i4/p833_1. See erratum [CAFE78].

Chasman:1978:ESS

[CAFE78] R. R. Chasman, I. Ahmad, A. M. Friedman, and J. R. Erskine. Erratum: Survey of single particle states in the mass region A > 228. Reviews of Modern Physics, 50(1):173, January 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.173;   http://rmp.aps.org/abstract/RMP/
v50/i1/p173_1. See [CAFE77].

Cowley:1972:PDM

[CB72] R. A. Cowley and W. J. L. Buyers. The properties of defects in magnetic insulators. Reviews of Modern Physics, 44 (2):406–450, April 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.406;   http://
rmp.aps.org/abstract/RMP/v44/i2/p406_1.

Chaloupka:1975:RPP

[CBBG^{+}75] Vladimir Chaloupka, Claude Bricman, Angela Barbaro-Galtieri, Denyse M. Chew, Robert L. Kelly, Thomas A. Lasinski, Alan Rittenberg, Arthur H. Rosenfeld, Thomas G. Trippe, Fumiyo Uchiyama, George P. Yost, Naomi Barash-Schmidt, and Matts Roos (Particle Data Group). Review of particle properties: Supplement to 1974 edition. Reviews of Modern Physics, 47(2):535–541, April 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.535;   http://
rmp.aps.org/abstract/RMP/v47/i2/p535_1.

Crispin:1970:DEI

[CF70] A. Crispin and G. N. Fowler. Density effect in the ionization energy loss of fast charged particles in matter. Reviews of Modern Physics, 42(3):290–316, July 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.42.290;   http:
//rmp.aps.org/abstract/RMP/v42/i3/p290_1.

Calmet:1977:AMM

[CNPdR77] J. Calmet, S. Narison, M. Perrottet, and E. de Rafael. The anomalous magnetic moment of the muon: a review of the theoretical contributions. Reviews of Modern Physics, 49(1):21–29, January 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.49.21;   http://rmp.aps.org/abstract/
RMP/v49/i1/p21_1.

Corwin:1975:GLA

[CNS75] L. Corwin, Y. Ne’eman, and S. Sternberg. Graded Lie algebras in mathematics and physics (bose–fermi symmetry). Reviews of Modern Physics, 47(3):573–603, July 1975. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.573;   http://rmp.aps.org/abstract/RMP/
v47/i3/p573_1.

Cohen:1977:HLR

[Coh77] Bernard L. Cohen. High-level radioactive waste from light-water reactors. Reviews of Modern Physics, 49(1):1–20, January 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.1;   http://rmp.aps.org/abstract/RMP/v49/
i1/p1_1.

Cole:1974:ESS

[Col74] Milton W. Cole. Electronic surface states of liquid helium. Reviews of Modern Physics, 46(3):451–464, July 1974. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.451;   http://rmp.aps.org/abstract/RMP/
v46/i3/p451_1.

Creutz:1978:FRL

[Cre78] Michael Creutz. Feynman rules for lattice gauge theory. Reviews of Modern Physics, 50(3):561–571, July 1978. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.561;   http://rmp.aps.org/abstract/RMP/
v50/i3/p561_1.

Clayton:1974:TA

[CW74] Donald D. Clayton and Stanford E. Woosley. Thermonuclear astrophysics. Reviews of Modern Physics, 46(4):755– 771, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.755;   http://
rmp.aps.org/abstract/RMP/v46/i4/p755_1.

Davidson:1972:PUN

[Dav72] Ernest R. Davidson. Properties and uses of natural orbitals. Reviews of Modern Physics, 44(3):451–464, July 1972. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.451;   http://rmp.aps.org/abstract/RMP/
v44/i3/p451_1.

Day:1978:CSN

[Day78] B. D. Day. Current state of nuclear matter calculations. Reviews of Modern Physics, 50(3):495–521, July 1978. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.495;   http://rmp.aps.org/abstract/RMP/
v50/i3/p495_1.

Devlin:1979:WHD

[DD79] Thomas J. Devlin and Jean O. Dickey. Weak hadronic decays: K → 2π and K → 3π. Reviews of Modern Physics, 51 (1):237–250, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.237;   http://
rmp.aps.org/abstract/RMP/v51/i1/p237_1.

Dean:1972:VPD

[Dea72] P. Dean. The vibrational properties of disordered systems: Numerical studies. Reviews of Modern Physics, 44(2):127– 168, April 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.127;   http://
rmp.aps.org/abstract/RMP/v44/i2/p127_1.

Deb:1973:FCC

[Deb73] B. M. Deb. The force concept in chemistry. Reviews of Modern Physics, 45(1):22–43, January 1973. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
45.22;   http://rmp.aps.org/abstract/RMP/v45/i1/p22_1.

Duvall:1977:PTU

[DG77] G. E. Duvall and R. A. Graham. Phase transitions under shock-wave loading. Reviews of Modern Physics, 49(3): 523–579, July 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.523;   http://
rmp.aps.org/abstract/RMP/v49/i3/p523_1.

Davidson:1979:ELQ

[DN79] Kris Davidson and Hagai Netzer. The emission lines of quasars and similar objects. Reviews of Modern Physics, 51 (4):715–766, October 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.715;   http://
rmp.aps.org/abstract/RMP/v51/i4/p715_1.

Dyson:1979:TEP

[Dys79] Freeman J. Dyson. Time without end: Physics and biology in an open universe. Reviews of Modern Physics, 51 (3):447–460, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.447;   http://
rmp.aps.org/abstract/RMP/v51/i3/p447_1.

Eden:1971:THE

[Ede71] R. J. Eden. Theorems on high energy collisions of elementary particles. Reviews of Modern Physics, 43(1):15–35, January 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.43.15;   http://rmp.aps.org/abstract/
RMP/v43/i1/p15_1.

Essam:1970:SBD

[EF70] John W. Essam and Michael E. Fisher. Some basic definitions in graph theory. Reviews of Modern Physics, 42 (2):271–288, April 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.42.271;   http://
rmp.aps.org/abstract/RMP/v42/i2/p271_1.

Elliott:1974:TPR

[EKL74] R. J. Elliott, J. A. Krumhansl, and P. L. Leath. The theory and properties of randomly disordered crystals and related physical systems. Reviews of Modern Physics, 46(3):465–543, July 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.465;   http://rmp.aps.org/abstract/RMP/
v46/i3/p465_1.

Enz:1974:TFH

[Enz74] Charles P. Enz. Two-fluid hydrodynamic description of ordered systems. Reviews of Modern Physics, 46(4):705– 753, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.705;   http://
rmp.aps.org/abstract/RMP/v46/i4/p705_1.

Ellis:1977:IGE

[EO77] P. J. Ellis and E. Osnes. An introductory guide to effective operators in nuclei. Reviews of Modern Physics, 49(4): 777–832, October 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.777;   http://
rmp.aps.org/abstract/RMP/v49/i4/p777_1.

Ellis:1977:LPD

[ES77] S. D. Ellis and R. Stroynowski. Large p_{T} physics: Data and the constituent models. Reviews of Modern Physics, 49 (4):753–775, October 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.753;   http://
rmp.aps.org/abstract/RMP/v49/i4/p753_1.

Esaki:1974:LJT

[Esa74] Leo Esaki. Long journey into tunneling. Reviews of Modern Physics, 46(2):237–244, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.237;   http:
//rmp.aps.org/abstract/RMP/v46/i2/p237_1.

Fagg:1975:ENM

[Fag75] Lawrence W. Fagg. Electroexcitation of nuclear magnetic dipole transitions. Reviews of Modern Physics, 47(3):683–711, July 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.683;   http://rmp.aps.org/abstract/RMP/
v47/i3/p683_1.

Fall:1979:GCC

[Fal79] S. Michael Fall. Galaxy correlations and cosmology. Reviews of Modern Physics, 51(1):21–42, January 1979. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
51.21;   http://rmp.aps.org/abstract/RMP/v51/i1/p21_1.

Feshbach:1974:DSP

[Fes74] Herman Feshbach. Doorway states and pre-equilibrium reactions. Reviews of Modern Physics, 46(1):1–5, January 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
46.1;   http://rmp.aps.org/abstract/RMP/v46/i1/p1_1.

Frazer:1972:HEM

[FIM^{+}72] W. R. Frazer, L. Ingber, C. H. Mehta, C. H. Poon, D. Silverman, K. Stowe, P. D. Ting, and H. J. Yesian. High-energy multiparticle reactions. Reviews of Modern Physics, 44(2):284–319, April 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.284;   http://rmp.aps.org/abstract/RMP/
v44/i2/p284_1.

Fisher:1974:RGT

[Fis74] Michael E. Fisher. The renormalization group in the theory of critical behavior. Reviews of Modern Physics, 46(4): 597–616, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.597;   http:
//rmp.aps.org/abstract/RMP/v46/i4/p597_1. See erratum

[Fis75].

Fisher:1975:ERG

[Fis75] Michael E. Fisher. Erratum: The renormalization group in the theory of critical behavior. Reviews of Modern Physics, 47(2):543, April 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.543;   http://
rmp.aps.org/abstract/RMP/v47/i2/p543_1. See [Fis74].

Frank:1971:SP

[FLS71] William M. Frank, David J. Land, and Richard M. Spector. Singular potentials. Reviews of Modern Physics, 43(1):36–98, January 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.43.36;   http://rmp.aps.org/abstract/
RMP/v43/i1/p36_1.

Fano:1973:IEP

[FM73] U. Fano and Joseph H. Macek. Impact excitation and polarization of the emitted light. Reviews of Modern Physics, 45 (4):553–573, October 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.45.553;   http://
rmp.aps.org/abstract/RMP/v45/i4/p553_1.

Fano:1976:EOP

[FTD76] U. Fano, C. E. Theodosiou, and J. L. Dehmer. Electronoptical properties of atomic fields. Reviews of Modern Physics, 48(1):49–68, January 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.48.49;   http://
rmp.aps.org/abstract/RMP/v48/i1/p49_1.

Gemmell:1974:CRE

[Gem74] Donald S. Gemmell. Channeling and related effects in the motion of charged particles through crystals. Reviews of Modern Physics, 46(1):129–227, January 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.129;   http:
//rmp.aps.org/abstract/RMP/v46/i1/p129_1.

Garcia:1973:ISV

[GFK73] J. D. Garcia, R. J. Fortner, and T. M. Kavanagh. Inner-shell vacancy production in ion-atom collisions. Reviews of Modern Physics, 45(2):111–177, April 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.45.111;   http:
//rmp.aps.org/abstract/RMP/v45/i2/p111_1.

Giaever:1974:ETS

[Gia74] Ivar Giaever. Electron tunneling and superconductivity. Reviews of Modern Physics, 46(2):245–250, April 1974. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.245;   http://rmp.aps.org/abstract/RMP/
v46/i2/p245_1.

Gitterman:1978:HFN

[Git78] Moshe Gitterman. Hydrodynamics of fluids near a critical point. Reviews of Modern Physics, 50(1):85–106, January 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
50.85;   http://rmp.aps.org/abstract/RMP/v50/i1/p85_1.

Griffin:1976:CHI

[GK76] James J. Griffin and Kit-Keung Kan. Colliding heavy ions: Nuclei as dynamical fluids. Reviews of Modern Physics, 48 (3):467–477, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.467;   http://
rmp.aps.org/abstract/RMP/v48/i3/p467_1.

Glendenning:1975:ICQ

[Gle75] Norman K. Glendenning. Interplay of classical and quantal descriptions of heavy-ion interactions. Reviews of Modern Physics, 47(3):659–682, July 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.659;   http://
rmp.aps.org/abstract/RMP/v47/i3/p659_1.

Gaillard:1975:SC

[GLR75] Mary K. Gaillard, Benjamin W. Lee, and Jonathan L. Rosner. Search for charm. Reviews of Modern Physics, 47 (2):277–310, April 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.277;   http://
rmp.aps.org/abstract/RMP/v47/i2/p277_1.

Golden:1971:LEE

[GLTG71] D. E. Golden, Neal F. Lane, A. Temkin, and E. Gerjuoy. Low energy electron-molecule scattering experiments and the theory of rotational excitation. Reviews of Modern Physics, 43 (4):642–678, October 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.43.642;   http://
rmp.aps.org/abstract/RMP/v43/i4/p642_1.

Goel:1971:VON

[GMM71] Narendra S. Goel, Samaresh C. Maitra, and Elliott W. Montroll. On the Volterra and other nonlinear models of interacting populations. Reviews of Modern Physics, 43(2):231–276, April 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.43.231;   http://rmp.aps.org/abstract/RMP/
v43/i2/p231_1.

Gell-Mann:1978:CEC

[GMRS78] M. Gell-Mann, P. Ramond, and R. Slansky. Color embeddings, charge assignments, and proton stability in unified gauge theories. Reviews of Modern Physics, 50(4):721–744, October 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.721;   http://rmp.aps.org/abstract/RMP/
v50/i4/p721_1.

Goldhaber:1971:TEL

[GN71] Alfred S. Goldhaber and Michael Martin Nieto. Terrestrial and extraterrestrial limits on the photon mass. Reviews of Modern Physics, 43(3):277–296, July 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.43.277;   http:
//rmp.aps.org/abstract/RMP/v43/i3/p277_1.

Greenberger:1979:CEN

[GO79] Daniel M. Greenberger and A. W. Overhauser. Coherence effects in neutron diffraction and gravity experiments. Reviews of Modern Physics, 51(1):43–78, January 1979. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
51.43;   http://rmp.aps.org/abstract/RMP/v51/i1/p43_1.

Gadzuk:1973:FEE

[GP73] J. W. Gadzuk and E. W. Plummer. Field Emission Energy Distribution (FEED). Reviews of Modern Physics, 45 (3):487–548, July 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.45.487;   http://
rmp.aps.org/abstract/RMP/v45/i3/p487_1.

Ginzburg:1976:EOC

[GP76a] V. L. Ginzburg and V. S. Ptuskin. Erratum: On the origin of cosmic rays: Some problems in high-energy astrophysics. Reviews of Modern Physics, 48(4):675–676, October 1976. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.675;   http://rmp.aps.org/abstract/RMP/
v48/i4/p675_1. See [GP76b].

Ginzburg:1976:OCR

[GP76b] V. L. Ginzburg and V. S. Ptuskin. On the origin of cosmic rays: Some problems in high-energy astrophysics. Reviews of Modern Physics, 48(2):161–189, April 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.48.161;   http:
//rmp.aps.org/abstract/RMP/v48/i2/p161_1. See erratum

[GP76a].

Gaisser:1978:CRS

[GPTM78] T. K. Gaisser, R. J. Protheroe, K. E. Turver, and T. J. L. McComb. Cosmic ray showers and particle physics at energies 10^{15}–10^{18} eV. Reviews of Modern Physics, 50(4): 859–880, October 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.859;   http:
//rmp.aps.org/abstract/RMP/v50/i4/p859_1. See erratum

[GPTM79].

Gaisser:1979:ECR

[GPTM79] T. K. Gaisser, R. J. Protheroe, K. E. Turver, and J. L. McComb. Erratum: Cosmic ray showers and particle physics at energies 10^{15}–10^{18} eV. Reviews of Modern Physics, 51(3):649, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.649;   http://rmp.aps.org/abstract/RMP/
v51/i3/p649_1. See [GPTM79].

Guyer:1971:EQC

[GRZ71] R. A. Guyer, R. C. Richardson, and L. I. Zane. Excitations in quantum crystals (A survey of NMR experiments in solid helium). Reviews of Modern Physics, 43(4):532–600, October 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.43.532;   http://rmp.aps.org/abstract/RMP/
v43/i4/p532_1.

Haken:1975:CPS

[Hak75] H. Haken. Cooperative phenomena in systems far from thermal equilibrium and in nonphysical systems. Reviews of Modern Physics, 47(1):67–121, January 1975. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
47.67;   http://rmp.aps.org/abstract/RMP/v47/i1/p67_1.

Happer:1972:OP

[Hap72] William Happer. Optical pumping. Reviews of Modern Physics, 44(2):169–249, April 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.169;   http://
rmp.aps.org/abstract/RMP/v44/i2/p169_1.

Haus:1979:PII

[Hau79] H. A. Haus. Physical interpretation of inverse scattering formalism applied to self-induced transparency. Reviews of Modern Physics, 51(2):331–339, April 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.51.331;   http:
//rmp.aps.org/abstract/RMP/v51/i2/p331_1.

Huebener:1974:MFS

[HC74] R. P. Huebener and John R. Clem. Magnetic flux structures in superconductors — a conference summary. Reviews of Modern Physics, 46(2):409–422, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.409;   http:
//rmp.aps.org/abstract/RMP/v46/i2/p409_1.

Hebel:1978:RAP

[HCD^{+}78] L. Charles Hebel, Eldon L. Christensen, Fred A. Donath, Warren E. Falconer, Leon J. Lidofsky, Ernest J. Moniz, Thomas H. Moss, Robert L. Pigford, Thomas H. Pigford, Gene I. Rochlin, Robert H. Silsbee, Mcdonald E. Wrenn, Hans Frauenfelder, Theodore L. Cairns, W. K. H. Panofsky, and M. Gene Simmons (APS Study Group Participants and APS Council Review Committee). Report to the American Physical Society by the study group on nuclear fuel cycles and waste management. Reviews of Modern Physics, 50(1):S1–S176, January 1978. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
50.S1;   http://rmp.aps.org/abstract/RMP/v50/i1/pS1_1.

Hewish:1975:PHD

[Hew75] A. Hewish. Pulsars and high density physics. Reviews of Modern Physics, 47(3):567–572, July 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.47.567;   http:
//rmp.aps.org/abstract/RMP/v47/i3/p567_1.

Hinton:1976:TPT

[HH76] F. L. Hinton and R. D. Hazeltine. Theory of plasma transport in toroidal confinement systems. Reviews of Modern Physics, 48(2):239–308, April 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.239;   http://
rmp.aps.org/abstract/RMP/v48/i2/p239_1.

Hohenberg:1977:TDC

[HH77] P. C. Hohenberg and B. I. Halperin. Theory of dynamic critical phenomena. Reviews of Modern Physics, 49(3):435– 479, July 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.435;   http://
rmp.aps.org/abstract/RMP/v49/i3/p435_1.

Hahn:1977:PPI

[HMR77] H. Hahn, M. Month, and R. R. Rau. Proton-proton intersecting storage accelerator facility ISABELLE at the Brookhaven National Laboratory. Reviews of Modern Physics, 49(3): 625–679, July 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.625;   http://
rmp.aps.org/abstract/RMP/v49/i3/p625_1.

Holstein:1974:REA

[Hol74] Barry R. Holstein. Recoil effects in allowed beta decay: The elementary particle approach. Reviews of Modern Physics, 46 (4):789–814, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.789;   http:
//rmp.aps.org/abstract/RMP/v46/i4/p789_1. See erratum

[Hol76].

Holstein:1976:ERE

[Hol76] Barry R. Holstein. Erratum: Recoil effects in allowed beta decay: The elementary particle approach. Reviews of Modern Physics, 48(4):673, October 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.48.673;   http:
//rmp.aps.org/abstract/RMP/v48/i4/p673_1. See [Hol74].

Harter:1978:FTR

[HPdP78] William G. Harter, Chris W. Patterson, and Fernando J. da Paixao. Frame transformation relations and multipole transitions in symmetric polyatomic molecules. Reviews of Modern Physics, 50(1):37–83, January 1978. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
50.37;   http://rmp.aps.org/abstract/RMP/v50/i1/p37_1.

Hurst:1979:RIS

[HPKY79] G. S. Hurst, M. G. Payne, S. D. Kramer, and J. P. Young. Resonance ionization spectroscopy and one-atom detection. Reviews of Modern Physics, 51(4):767–819, October 1979. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.767;   http://rmp.aps.org/abstract/RMP/
v51/i4/p767_1.

Huang:1979:GER

[Hua79] Keh-Ning Huang. Graphical evaluation of relativistic matrix elements. Reviews of Modern Physics, 51(1):215–236, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.215;   http://rmp.aps.org/abstract/RMP/
v51/i1/p215_1.

Hehl:1976:GRS

[HvdHKN76] Friedrich W. Hehl, Paul von der Heyde, G. David Kerlick, and James M. Nester. General Relativity with spin and torsion: Foundations and prospects. Reviews of Modern Physics, 48(3):393–416, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.393;   http://
rmp.aps.org/abstract/RMP/v48/i3/p393_1.

Igo:1978:SRI

[Igo78] G. J. Igo. Some recent intermediate-and high-energy protonnucleus research. Reviews of Modern Physics, 50(3):523– 560, July 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.50.523;   http://
rmp.aps.org/abstract/RMP/v50/i3/p523_1.

Iliopoulos:1975:FMP

[IIM75] J. Iliopoulos, C. Itzykson, and A. Martin. Functional methods and perturbation theory. Reviews of Modern Physics, 47 (1):165–192, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.165;   http://
rmp.aps.org/abstract/RMP/v47/i1/p165_1.

Inokuti:1978:AIC

[IIT78] Mitio Inokuti, Yukikazu Itikawa, and James E. Turner. Addenda: Inelastic collisions of fast charged particles with atoms and molecules — the Bethe theory revisited. Reviews of Modern Physics, 50(1):23–35, January 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.23;   http:/
/rmp.aps.org/abstract/RMP/v50/i1/p23_1. See [Ino71].

Inglis:1975:TDS

[Ing75] David Rittenhouse Inglis. The two-dimensionality of slow motion in a rotating fluid. Reviews of Modern Physics, 47(4): 841–847, October 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.841;   http://
rmp.aps.org/abstract/RMP/v47/i4/p841_1.

Inokuti:1971:ICF

[Ino71] Mitio Inokuti. Inelastic collisions of fast charged particles with atoms and molecules — the Bethe theory revisited. Reviews of Modern Physics, 43(3):297–347, July 1971. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.43.297;   http://rmp.aps.org/abstract/RMP/

v43/i3/p297_1. See addendum [IIT78].

Jackson:1970:MHE

[Jac70] J. D. Jackson. Models for high-energy processes. Reviews of Modern Physics, 42(1):12–67, January 1970. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
42.12;   http://rmp.aps.org/abstract/RMP/v42/i1/p12_1.

Jackson:1976:USF

[Jac76] J. D. Jackson. On understanding spin-flip synchrotron radiation and the transverse polarization of electrons in storage rings. Reviews of Modern Physics, 48(3):417–433, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.417;   http://rmp.aps.org/abstract/RMP/
v48/i3/p417_1.

Jackiw:1977:QMC

[Jac77] R. Jackiw. Quantum meaning of classical field theory. Reviews of Modern Physics, 49(3):681–706, July 1977. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.681;   http://rmp.aps.org/abstract/RMP/
v49/i3/p681_1.

Jacob:1973:QFS

[JM73] Gerhard Jacob and Th. A. J. Maris. Quasi-free scattering and nuclear structure. II. Reviews of Modern Physics, 45 (1):6–21, January 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.45.6;   http://
rmp.aps.org/abstract/RMP/v45/i1/p6_1.

Jones:1976:OGR

[Jon76] Bernard J. T. Jones. The origin of galaxies: a review of recent theoretical developments and their confrontation with observation. Reviews of Modern Physics, 48(1):107–149, January 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.107;   http://rmp.aps.org/abstract/RMP/
v48/i1/p107_1.

Jones:1977:RQS

[Jon77] Lawrence W. Jones. A review of quark search experiments. Reviews of Modern Physics, 49(4):717–752, October 1977. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.49.717;   http://rmp.aps.org/abstract/RMP/
v49/i4/p717_1.

Josephson:1974:DTS

[Jos74] B. D. Josephson. The discovery of tunnelling supercurrents. Reviews of Modern Physics, 46(2):251–254, April 1974. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.251;   http://rmp.aps.org/abstract/RMP/
v46/i2/p251_1.

Joachain:1974:MSE

[JQ74] Charles J. Joachain and C. Quigg. Multiple scattering expansions in several particle dynamics. Reviews of Modern Physics, 46(2):279–324, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.279;   http://
rmp.aps.org/abstract/RMP/v46/i2/p279_1.

Kadanoff:1977:ARG

[Kad77] Leo P. Kadanoff. The application of renormalization group techniques to quarks and strings. Reviews of Modern Physics, 49(2):267–296, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.267;   http://
rmp.aps.org/abstract/RMP/v49/i2/p267_1.

Kapitza:1979:PCT

[Kap79] P. L. Kapitza. Plasma and the controlled thermonuclear reaction. Reviews of Modern Physics, 51(3):417–423, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.417;   http://rmp.aps.org/abstract/RMP/
v51/i3/p417_1.

Kirkpatrick:1973:PC

[Kir73] Scott Kirkpatrick. Percolation and conduction. Reviews of Modern Physics, 45(4):574–588, October 1973. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.574;   http://rmp.aps.org/abstract/RMP/
v45/i4/p574_1.

Ma:1973:IRG

[kM73] Shang keng Ma. Introduction to the renormalization group. Reviews of Modern Physics, 45(4):589–614, October 1973. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.589;   http://rmp.aps.org/abstract/RMP/
v45/i4/p589_1.

Kogut:1979:ILG

[Kog79] John B. Kogut. An introduction to lattice gauge theory and spin systems. Reviews of Modern Physics, 51(4):659– 713, October 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.659;   http://
rmp.aps.org/abstract/RMP/v51/i4/p659_1.

Kaup:1979:EST

[KRB79a] D. J. Kaup, A. Reiman, and A. Bers. Erratum: Space–time evolution of nonlinear three-wave interactions. I. Interaction in a homogeneous medium. Reviews of Modern Physics, 51 (4):915–917, October 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.915;   http://
rmp.aps.org/abstract/RMP/v51/i4/p915_1. See [KRB79b].

Kaup:1979:STE

[KRB79b] D. J. Kaup, A. Reiman, and A. Bers. Space–time evolution of nonlinear three-wave interactions. I. Interaction in a homogeneous medium. Reviews of Modern Physics, 51(2):275–309, April 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.275;   http://rmp.aps.org/abstract/RMP/

v51/i2/p275_1. See erratum [KRB79a].

Kohn:1970:TKB

[KS70] W. Kohn and D. Sherrington. Two kinds of bosons and Bose condensates. Reviews of Modern Physics, 42(1):1–11, January 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.42.1;   http://rmp.aps.org/abstract/RMP/v42/
i1/p1_1.

Kane:1976:ITB

[KS76] G. L. Kane and A. Seidl. An interpretation of two-body hadron reactions. Reviews of Modern Physics, 48(2):309–355, April 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.309;   http://rmp.aps.org/abstract/RMP/
v48/i2/p309_1.

Kaufmann:1979:EFG

[KV79] Elton N. Kaufmann and Reiner J. Vianden. The electric field gradient in noncubic metals. Reviews of Modern Physics, 51 (1):161–214, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.161;   http://
rmp.aps.org/abstract/RMP/v51/i1/p161_1.

Lamb:1971:ADU

[Lam71] G. L. Lamb, Jr. Analytical descriptions of ultrashort optical pulse propagation in a resonant medium. Reviews of Modern Physics, 43(2):99–124, April 1971. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
43.99;   http://rmp.aps.org/abstract/RMP/v43/i2/p99_1.

Lewis:1975:RAP

[LBC^{+}75] H. W. Lewis, R. J. Budnitz, A. W. Castleman, D. E. Dorfan, F. C. Finlayson, R. L. Garwin, L. C. Hebel, S. M. Keeny, Jr., R. A. Muller, T. B. Taylor, G. F. Smoot, F. von Hippel, Hans Bethe, W. K. H. Panofsky, and V. F. Weisskopf (APS Study Group Participants and APS Council Review Committee). Report to the American Physical Society by the study group on light-water reactor safety. Reviews of Modern Physics, 47(S1):S1–S123, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.47.S1;   http://rmp.aps.org/abstract/
RMP/v47/iS1/pS1_1.

Lasinski:1973:RPP

[LBGK^{+}73] Thomas A. Lasinski, Angela Barbaro-Galtieri, Robert L. Kelly, Alan Rittenberg, Arthur H. Rosenfeld, Thomas G. Trippe, Naomi Barash-Schmidt, Claude Bricman, Vladimir Chaloupka, Paul Söding, and Matts Roos. Review of particle properties particle data group. Reviews of Modern Physics, 45(2):S1–S175, April 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.45.S1;   http://rmp.aps.org/abstract/
RMP/v45/i2/pS1_1.

Lee:1975:ANS

[Lee75] T. D. Lee. Abnormal nuclear states and vacuum excitation. Reviews of Modern Physics, 47(2):267–275, April 1975. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.267;   http://rmp.aps.org/abstract/RMP/
v47/i2/p267_1.

Leggett:1975:TDN

[Leg75] Anthony J. Leggett. A theoretical description of the new phases of liquid^{3}He. Reviews of Modern Physics, 47(2):331–414, April 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.331;   http://rmp.aps.org/abstract/RMP/

v47/i2/p331_1. See erratum [Leg76].

Leggett:1976:ETD

[Leg76] A. J. Leggett. Erratum: A theoretical description of the new phases of liquid ^{3}He. Reviews of Modern Physics, 48 (2):357, April 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.357;   http://
rmp.aps.org/abstract/RMP/v48/i2/p357_1. See [Leg75].

Leibbrandt:1975:ITD

[Lei75] George Leibbrandt. Introduction to the technique of dimensional regularization. Reviews of Modern Physics, 47(4): 849–876, October 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.849;   http://
rmp.aps.org/abstract/RMP/v47/i4/p849_1.

Langhoff:1972:ATD

[LEK72] P. W. Langhoff, S. T. Epstein, and M. Karplus. Aspects of time-dependent perturbation theory. Reviews of Modern Physics, 44(3):602–644, July 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.602;   http://
rmp.aps.org/abstract/RMP/v44/i3/p602_1.

Louck:1972:AOU

[LG72] James D. Louck and Harold W. Galbraith. Application of orthogonal and unitary group methods to the N-body problem. Reviews of Modern Physics, 44(3):540–601, July 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.540;   http://rmp.aps.org/abstract/RMP/
v44/i3/p540_1.

Louck:1976:EVE

[LG76] James D. Louck and Harold W. Galbraith. Eckart vectors, Eckart frames, and polyatomic molecules. Reviews of Modern Physics, 48(1):69–106, January 1976. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
48.69;   http://rmp.aps.org/abstract/RMP/v48/i1/p69_1.

Lifshitz:1978:SPS

[LGK78] I. M. Lifshitz, A. Yu. Grosberg, and A. R. Khokhlov. Some problems of the statistical physics of polymer chains with volume interaction. Reviews of Modern Physics, 50(3):683–713, July 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.683;   http://rmp.aps.org/abstract/RMP/
v50/i3/p683_1.

Lieb:1976:SM

[Lie76] Elliott H. Lieb. The stability of matter. Reviews of Modern Physics, 48(4):553–569, October 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.48.553;   http:
//rmp.aps.org/abstract/RMP/v48/i4/p553_1.

Likharev:1979:SWL

[Lik79] K. K. Likharev. Superconducting weak links. Reviews of Modern Physics, 51(1):101–159, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.51.101;   http:
//rmp.aps.org/abstract/RMP/v51/i1/p101_1.

Laubereau:1978:VDL

[LK78] A. Laubereau and W. Kaiser. Vibrational dynamics of liquids and solids investigated by picosecond light pulses. Reviews of Modern Physics, 50(3):607–665, July 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.607;   http:
//rmp.aps.org/abstract/RMP/v50/i3/p607_1.

Lightman:1978:DEG

[LS78] Alan P. Lightman and Stuart L. Shapiro. The dynamical evolution of globular clusters. Reviews of Modern Physics, 50 (2):437–481, April 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.50.437;   http://
rmp.aps.org/abstract/RMP/v50/i2/p437_1.

Maris:1977:PPI

[Mar77] Humphrey J. Maris. Phonon-phonon interactions in liquid helium. Reviews of Modern Physics, 49(2):341–359, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.341;   http://rmp.aps.org/abstract/RMP/
v49/i2/p341_1.

McRae:1979:ESR

[McR79] E. G. McRae. Electronic surface resonances of crystals. Reviews of Modern Physics, 51(3):541–568, July 1979. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.541;   http://rmp.aps.org/abstract/RMP/
v51/i3/p541_1.

Mermin:1979:TTD

[Mer79] N. D. Mermin. The topological theory of defects in ordered media. Reviews of Modern Physics, 51(3):591–648, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.591;   http://rmp.aps.org/abstract/RMP/
v51/i3/p591_1.

Migdal:1978:PFN

[Mig78] A. B. Migdal. Pion fields in nuclear matter. Reviews of Modern Physics, 50(1):107–172, January 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.107;   http:
//rmp.aps.org/abstract/RMP/v50/i1/p107_1.

Morse:1975:EUC

[Mor75] Philip M. Morse. Edward Uhler Condon, 1902–1974. Reviews of Modern Physics, 47(1):1–6, January 1975. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
47.1;   http://rmp.aps.org/abstract/RMP/v47/i1/p1_1.

Mottelson:1976:EME

[Mot76] Ben Mottelson. Elementary modes of excitation in the nucleus. Reviews of Modern Physics, 48(3):375–383, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.375;   http://rmp.aps.org/abstract/RMP/
v48/i3/p375_1.

Mott:1978:EG

[Mot78] Sir Nevill Mott. Electrons in glass. Reviews of Modern Physics, 50(2):203–208, April 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.50.203;   http://
rmp.aps.org/abstract/RMP/v50/i2/p203_1.

Moldover:1979:GEF

[MSGH79] M. R. Moldover, J. V. Sengers, R. W. Gammon, and R. J. Hocken. Gravity effects in fluids near the gas-liquid critical point. Reviews of Modern Physics, 51(1):79–99, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
51.79;   http://rmp.aps.org/abstract/RMP/v51/i1/p79_1.

Narayanamurti:1970:TSD

[NP70] V. Narayanamurti and R. O. Pohl. Tunneling states of defects in solids. Reviews of Modern Physics, 42(2):201–236, April 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.42.201;   http://rmp.aps.org/abstract/RMP/
v42/i2/p201_1.

Normand:1977:CIP

[NPV77] Christiane Normand, Yves Pomeau, and Manuel G. Velarde. Convective instability: a physicist’s approach. Reviews of Modern Physics, 49(3):581–624, July 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.49.581;   http:
//rmp.aps.org/abstract/RMP/v49/i3/p581_1.

Ogurtsov:1972:ESE

[Ogu72a] Gennadi N. Ogurtsov. Energy spectra of electrons ejected in ion-atom collisions. Reviews of Modern Physics, 44(1): 1–17, January 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.44.1;   http://
rmp.aps.org/abstract/RMP/v44/i1/p1_1. See erratum [Ogu72b].

Ogurtsov:1972:EES

[Ogu72b] Gennadi N. Ogurtsov. Erratum: Energy Spectra of Electrons Ejected in Ion-Atom Collisions. Reviews of Modern Physics, 44(4):814, October 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.814;   http://
rmp.aps.org/abstract/RMP/v44/i4/p814_1. See [Ogu72a].

Olshansky:1979:PGO

[Ols79] R. Olshansky. Propagation in glass optical waveguides. Reviews of Modern Physics, 51(2):341–367, April 1979. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.341;   http://rmp.aps.org/abstract/RMP/
v51/i2/p341_1.

Ormonde:1975:EVR

[Orm75a] Stephan Ormonde. Erratum: Vibrational relaxation theories and measurements. Reviews of Modern Physics, 47(4): 977, October 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.977;   http://
rmp.aps.org/abstract/RMP/v47/i4/p977_1. See [Orm75b].

Ormonde:1975:VRT

[Orm75b] Stephan Ormonde. Vibrational relaxation theories and measurements. Reviews of Modern Physics, 47(1):193–258, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.193;   http://rmp.aps.org/abstract/RMP/

v47/i1/p193_1. See erratum [Orm75a].

Orzalesi:1970:CGS

[Orz70] Claudio A. Orzalesi. Charges and generators of symmetry transformations in quantum field theory. Reviews of Modern Physics, 42(4):381–408, October 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.42.381;   http://
rmp.aps.org/abstract/RMP/v42/i4/p381_1.

Ogle:1971:SPL

[OWPF71] W. Ogle, S. Wahlborn, R. Piepenbring, and S. Fredriksson. Singleparticle levels of nonspherical nuclei in the region 150 < A < 190. Reviews of Modern Physics, 43(3):424–478, July 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.43.424;   http://rmp.aps.org/abstract/RMP/
v43/i3/p424_1.

Pais:1977:RTE

[Pai77] A. Pais. Radioactivity’s two early puzzles. Reviews of Modern Physics, 49(4):925–938, October 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.49.925;   http:
//rmp.aps.org/abstract/RMP/v49/i4/p925_1.

Pais:1979:EQT

[Pai79] A. Pais. Einstein and the quantum theory. Reviews of Modern Physics, 51(4):863–914, October 1979. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/abstract/RMP/v51/p863;
 http://link.aps.org/doi/10.1103/RevModPhys.51.863;
http://prola.aps.org/abstract/RMP/v51/i4/p863_1;   http:
//prola.aps.org/pdf/RMP/v51/i4/p863_1;   http://rmp.aps.
org/abstract/RMP/v51/i4/p863_1.

Pantelides:1978:ESI

[Pan78] Sokrates T. Pantelides. The electronic structure of impurities and other point defects in semiconductors. Reviews of Modern Physics, 50(4):797–858, October 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.797;   http:
//rmp.aps.org/abstract/RMP/v50/i4/p797_1.

PDG:1976:ERP

[Par76] Particle Data Group. Erratum: Review of particle properties. Reviews of Modern Physics, 48(3):497, July 1976. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.497;   http://rmp.aps.org/abstract/RMP/

v48/i3/p497_1. See [TBGK^{+}76].

Porkolab:1978:NWE

[PC78] M. Porkolab and R. P. H. Chang. Nonlinear wave effects in laboratory plasmas: a comparison between theory and experiment. Reviews of Modern Physics, 50(4):745–795, October 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.745;   http://rmp.aps.org/abstract/RMP/
v50/i4/p745_1.

Pines:1976:ERD

[PCF^{+}76] David Pines, E. D. Commins, H. P. Furth, R. L. Garwin, B. I. Halperin, R. W. Keyes, E. E. Salpeter, J. P. Schiffer, S. B. Treiman, and G. K. Walters. Editorial: Reviews dealing with applications of physics. Reviews of Modern Physics, 48(4):511, October 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.511;   http://
rmp.aps.org/abstract/RMP/v48/i4/p511_1.

Penzias:1979:OE

[Pen79] Arno A. Penzias. The origin of the elements. Reviews of Modern Physics, 51(3):425–431, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.51.425;   http:
//rmp.aps.org/abstract/RMP/v51/i3/p425_1.

Pines:1974:ESN

[PGH^{+}74] David Pines, Roy W. Gould, Bertrand I. Halperin, Rolf Landauer, Arthur V. Phelps, Edwin E. Salpeter, John Schiffer, and Sam B. Treiman. Editorial: Some new directions. Reviews of Modern Physics, 46(2):235, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.235;   http://
rmp.aps.org/abstract/RMP/v46/i2/p235_1.

Phillips:1970:ICB

[Phi70] J. C. Phillips. Ionicity of the chemical bond in crystals. Reviews of Modern Physics, 42(3):317–356, July 1970. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.42.317;   http://rmp.aps.org/abstract/RMP/
v42/i3/p317_1.

Pines:1977:IOR

[Pin77] David Pines. An invitation to our readers. Reviews of Modern Physics, 49(2):233, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.233;   http://
rmp.aps.org/abstract/RMP/v49/i2/p233_1.

Powell:1976:CSI

[Pow76] C. J. Powell. Cross sections for ionization of inner-shell electrons by electrons. Reviews of Modern Physics, 48(1):33–47, January 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.48.33;   http://rmp.aps.org/abstract/
RMP/v48/i1/p33_1.

Pratt:1973:EAP

[PRT73a] R. H. Pratt, A. Ron, and H. K. Tseng. Erratum: Atomic Photoelectric Effect above 10 keV. Reviews of Modern Physics, 45 (4):663–664, October 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.45.663;   http://
rmp.aps.org/abstract/RMP/v45/i4/p663_1. See [PRT73b].

Pratt:1973:APE

[PRT73b] R. H. Pratt, Akiva Ron, and H. K. Tseng. Atomic photoelectric effect above 10 keV. Reviews of Modern Physics, 45 (2):273–325, April 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.45.273;   http:
//rmp.aps.org/abstract/RMP/v45/i2/p273_1. See erratum

[PRT73a].

Pandharipande:1979:VTN

[PW79] V. R. Pandharipande and R. B. Wiringa. Variations on a theme of nuclear matter. Reviews of Modern Physics, 51(4): 821–861, October 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.821;   http://
rmp.aps.org/abstract/RMP/v51/i4/p821_1.

Quigg:1977:PDI

[Qui77] C. Quigg. Production and detection of intermediate vector bosons and heavy leptons in pp and p̄p collisions. Reviews of Modern Physics, 49(2):297–315, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.49.297;   http:
//rmp.aps.org/abstract/RMP/v49/i2/p297_1.

Rainwater:1976:BSN

[Rai76] James Rainwater. Background for the spheroidal nuclear model proposal. Reviews of Modern Physics, 48(3):385–391, July 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.385;   http://rmp.aps.org/abstract/RMP/
v48/i3/p385_1.

Rittenberg:1971:RPP

[RBGL^{+}71] Alan Rittenberg, Angela Barbaro-Galtieri, Thomas Lasinski, Arthur H. Rosenfeld, Thomas G. Trippe, Matts Roos, Claude Bricman, Paul Söding, Naomi Barash-Schmidt, and Charles G. Wohl. Review of particle properties Particle Data Group. Reviews of Modern Physics, 43(2):S1–S150, April 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
43.S1;   http://rmp.aps.org/abstract/RMP/v43/i2/pS1_1.

Reiman:1979:STE

[Rei79] A. Reiman. Space–time evolution of nonlinear three-wave interactions. II. Interaction in an inhomogeneous medium. Reviews of Modern Physics, 51(2):311–330, April 1979. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.311;   http://rmp.aps.org/abstract/RMP/
v51/i2/p311_1.

Reya:1974:CSB

[Rey74] E. Reya. Chiral symmetry breaking and meson-nucleon sigma commutators: a review. Reviews of Modern Physics, 46 (3):545–580, July 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.545;   http://
rmp.aps.org/abstract/RMP/v46/i3/p545_1.

Rujula:1974:FFN

[RGGQ74] A. De. Rújula, Howard Georgi, S. L. Glashow, and Helen R. Quinn. Fact and fancy in neutrino physics. Reviews of Modern Physics, 46(2):391–407, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.391;   http:
//rmp.aps.org/abstract/RMP/v46/i2/p391_1.

Ribe:1975:FRS

[Rib75] F. L. Ribe. Fusion reactor systems. Reviews of Modern Physics, 47(1):7–41, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.47.7;   http://
rmp.aps.org/abstract/RMP/v47/i1/p7_1.

Richter:1977:PCE

[Ric77] Burton Richter. From the psi to charm: The experiments of 1975 and 1976. Reviews of Modern Physics, 49(2):251–266, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.251;   http://rmp.aps.org/abstract/RMP/
v49/i2/p251_1.

Rowe:1975:TEM

[RNT75] D. J. Rowe and C. Ngo-Trong. Tensor equations of motion for the excitations of rotationally invariant or charge-independent systems. Reviews of Modern Physics, 47(2):471–485, April 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.47.471;   http://rmp.aps.org/abstract/RMP/
v47/i2/p471_1.

Rebbi:1970:CMS

[RS70] Claudio Rebbi and Richard Slansky. Crossing matrices for SU(2) and SU(3). Reviews of Modern Physics, 42(1):68–86, January 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.42.68;   http://rmp.aps.org/abstract/
RMP/v42/i1/p68_1.

Rich:1972:CSL

[RW72] Arthur Rich and John C. Wesley. The current status of the lepton g factors. Reviews of Modern Physics, 44(2):250–283, April 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.250;   http://rmp.aps.org/abstract/RMP/
v44/i2/p250_1.

Ryle:1975:RTL

[Ryl75] M. Ryle. Radio telescopes of large resolving power. Reviews of Modern Physics, 47(3):557–566, July 1975. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.557;   http://rmp.aps.org/abstract/RMP/
v47/i3/p557_1.

Sagdeev:1979:OLCa

[Sag79a] Roald Z. Sagdeev. The 1976 Oppenheimer Lectures: Critical problems in plasma astrophysics. I. Turbulence and nonlinear waves. Reviews of Modern Physics, 51(1):1–9, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.1;   http://rmp.aps.org/abstract/RMP/v51/
i1/p1_1.

Sagdeev:1979:OLCb

[Sag79b] Roald Z. Sagdeev. The 1976 Oppenheimer Lectures: Critical problems in plasma astrophysics. II. Singular layers and reconnection. Reviews of Modern Physics, 51(1):11–20, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.51.11;   http://rmp.aps.org/abstract/
RMP/v51/i1/p11_1.

Salpeter:1974:DSR

[Sal74] E. E. Salpeter. Dying stars and reborn dust. Reviews of Modern Physics, 46(3):433–436, July 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.433;   http:
//rmp.aps.org/abstract/RMP/v46/i3/p433_1.

Salour:1978:QIE

[Sal78] M. M. Salour. Quantum interference effects in two-photon spectroscopy. Reviews of Modern Physics, 50(3):667–681, July 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.667;   http://rmp.aps.org/abstract/RMP/
v50/i3/p667_1.

Satchler:1978:ODT

[Sat78] G. Ray Satchler. Outline of the development of the theory of direct nuclear reactions. Reviews of Modern Physics, 50(1):1–10, January 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.1;   http://
rmp.aps.org/abstract/RMP/v50/i1/p1_1.

Schulz:1973:REIa

[Sch73a] George J. Schulz. Resonances in electron impact on atoms. Reviews of Modern Physics, 45(3):378–422, July 1973. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.378;   http://rmp.aps.org/abstract/RMP/
v45/i3/p378_1.

Schulz:1973:REIb

[Sch73b] George J. Schulz. Resonances in electron impact on diatomic molecules. Reviews of Modern Physics, 45(3):423–486, July 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.423;   http://rmp.aps.org/abstract/RMP/
v45/i3/p423_1.

Scherk:1975:ITD

[Sch75] J. Scherk. An introduction to the theory of dual models and strings. Reviews of Modern Physics, 47(1):123–164, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.123;   http://rmp.aps.org/abstract/RMP/
v47/i1/p123_1.

Schnakenberg:1976:NTM

[Sch76] J. Schnakenberg. Network theory of microscopic and macroscopic behavior of master equation systems. Reviews of Modern Physics, 48(4):571–585, October 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.48.571;   http://
rmp.aps.org/abstract/RMP/v48/i4/p571_1.

Scott:1974:SMS

[Sco74] J. F. Scott. Soft-mode spectroscopy: Experimental studies of structural phase transitions. Reviews of Modern Physics, 46 (1):83–128, January 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.46.83;   http://
rmp.aps.org/abstract/RMP/v46/i1/p83_1.

Scott:1975:ENF

[Sco75] Alwyn C. Scott. The electrophysics of a nerve fiber. Reviews of Modern Physics, 47(2):487–533, April 1975. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.487;   http://rmp.aps.org/abstract/RMP/
v47/i2/p487_1.

Samios:1974:HCR

[SGM74] N. P. Samios, M. Goldberg, and B. T. Meadows. Hadrons and SU(3): a critical review. Reviews of Modern Physics, 46 (1):49–81, January 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.46.49;   http://
rmp.aps.org/abstract/RMP/v46/i1/p49_1.

Shank:1975:PDL

[Sha75] C. V. Shank. Physics of dye lasers. Reviews of Modern Physics, 47(3):649–657, July 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.649;   http://
rmp.aps.org/abstract/RMP/v47/i3/p649_1.

Shen:1976:RAN

[She76] Y. R. Shen. Recent advances in nonlinear optics. Reviews of Modern Physics, 48(1):1–32, January 1976. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
48.1;   http://rmp.aps.org/abstract/RMP/v48/i1/p1_1.

Shirane:1974:NSS

[Shi74] G. Shirane. Neutron scattering studies of structural phase transitions at Brookhaven. Reviews of Modern Physics, 46 (3):437–449, July 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.437;   http://
rmp.aps.org/abstract/RMP/v46/i3/p437_1.

Sinfelt:1979:SMC

[Sin79] John H. Sinfelt. Structure of metal catalysts. Reviews of Modern Physics, 51(3):569–589, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.51.569;   http:
//rmp.aps.org/abstract/RMP/v51/i3/p569_1.

Sirlin:1978:CAF

[Sir78a] A. Sirlin. Current algebra formulation of radiative corrections in gauge theories and the universality of the weak interactions. Reviews of Modern Physics, 50(3):573–605, July 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.573;   http://rmp.aps.org/abstract/RMP/

v50/i3/p573_1. See erratum [Sir78b].

Sirlin:1978:ECA

[Sir78b] A. Sirlin. Erratum: Current algebra formulation of radiative corrections in gauge theories and the universality of the weak interactions. Reviews of Modern Physics, 50(4):905, October 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.50.905;   http://rmp.aps.org/abstract/RMP/
v50/i4/p905_1. See [Sir78a].

Sorensen:1973:NMI

[Sor73] Raymond A. Sorensen. Nuclear moment of inertia at high spin. Reviews of Modern Physics, 45(3):353–377, July 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.353;   http://rmp.aps.org/abstract/RMP/
v45/i3/p353_1.

Specht:1974:NF

[Spe74] Hans J. Specht. Nuclear fission. Reviews of Modern Physics, 46(4):773–787, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.46.773;   http://
rmp.aps.org/abstract/RMP/v46/i4/p773_1.

Stephen:1974:PLC

[SS74] Michael J. Stephen and Joseph P. Straley. Physics of liquid crystals. Reviews of Modern Physics, 46(4):617–704, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.617;   http://rmp.aps.org/abstract/RMP/
v46/i4/p617_1.

Shakeshaft:1979:MCT

[SS79] Robin Shakeshaft and Larry Spruch. Mechanisms for charge transfer (or for the capture of any light particle) at asymptotically high impact velocities. Reviews of Modern Physics, 51(2):369–405, April 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.51.369;   http://
rmp.aps.org/abstract/RMP/v51/i2/p369_1.

Schiffer:1976:EIB

[ST76] John P. Schiffer and William W. True. The effective interaction between nucleons deduced from nuclear spectra. Reviews of Modern Physics, 48(2):191–217, April 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.48.191;   http:
//rmp.aps.org/abstract/RMP/v48/i2/p191_1.

Stephens:1975:CER

[Ste75] F. S. Stephens. Coriolis effects and rotation alignment in nuclei. Reviews of Modern Physics, 47(1):43–65, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
47.43;   http://rmp.aps.org/abstract/RMP/v47/i1/p43_1.

Sivers:1971:RRW

[SY71] Dennis Sivers and Joel Yellin. Review of recent work on narrow resonance models. Reviews of Modern Physics, 43 (2):125–188, April 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.43.125;   http://
rmp.aps.org/abstract/RMP/v43/i2/p125_1.

Trippe:1976:RPP

[TBGK^{+}76] Thomas G. Trippe, Angela Barbaro-Galtieri, Robert L. Kelly, Alan Rittenberg, Arthur H. Rosenfeld, George P. Yost, Naomi Barash-Schmidt, Claude Bricman, Richard J. Hemingway, Michael J. Losty, Matts Roos, Vladimir Chaloupka, and Betty Armstrong (Particle Data Group). Review of particle properties. Reviews of Modern Physics, 48(2):S1–S245, April 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.48.S1;   http://rmp.aps.org/abstract/

RMP/v48/i2/pS1_1. See erratum [Par76].

Terazawa:1973:TPP

[Ter73] Hidezumi Terazawa. Two-photon processes for particle production at high energies. Reviews of Modern Physics, 45(4): 615–662, October 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.45.615;   http://
rmp.aps.org/abstract/RMP/v45/i4/p615_1.

Testardi:1975:SIS

[Tes75] L. R. Testardi. Structural instability and superconductivity in A-15 compounds. Reviews of Modern Physics, 47 (3):637–648, July 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.637;   http://
rmp.aps.org/abstract/RMP/v47/i3/p637_1.

Tien:1977:ION

[Tie77] P. K. Tien. Integrated optics and new wave phenomena in optical waveguides. Reviews of Modern Physics, 49(2):

361–420, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.361;   http://
rmp.aps.org/abstract/RMP/v49/i2/p361_1.

Tinkham:1974:EPS

[Tin74] M. Tinkham. The electromagnetic properties of superconductors. Reviews of Modern Physics, 46(4):587–596, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.46.587;   http://rmp.aps.org/abstract/RMP/
v46/i4/p587_1.

Ting:1977:DJP

[Tin77] Samuel C. C. Ting. The discovery of the J particle: a personal recollection. Reviews of Modern Physics, 49(2):235–249, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.235;   http://rmp.aps.org/abstract/RMP/
v49/i2/p235_1.

Trickey:1972:TEM

[TKA72] S. B. Trickey, W. P. Kirk, and E. D. Adams. Thermodynamic, elastic, and magnetic properties of solid helium. Reviews of Modern Physics, 44(4):668–715, October 1972. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.44.668;   http://rmp.aps.org/abstract/RMP/
v44/i4/p668_1.

Taylor:1969:DUM

[TPL69] B. N. Taylor, W. H. Parker, and D. N. Langenberg. Determination of e/h, using macroscopic quantum phase coherence in superconductors: Implications for quantum electrodynamics and the fundamental physical constants. Reviews of Modern Physics, 41(3):375–496, July 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.41.375;   http:
//rmp.aps.org/abstract/RMP/v41/i3/p375_1.  See errata

[TPL73].

Taylor:1973:EDU

[TPL73] B. N. Taylor, W. H. Parker, and D. N. Langenberg. Erratum: Determination of e/h, Using Macroscopic Quantum Phase Coherence in Superconductors: Implications for Quantum Electrodynamics and the Fundamental Physical Constants. Reviews of Modern Physics, 45(1):109, January 1973. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.109;   http://rmp.aps.org/abstract/RMP/
v45/i1/p109_1. See [TPL69].

Tawara:1973:CCP

[TR73a] H. Tawara and A. Russek. Charge changing processes in hydrogen beams. Reviews of Modern Physics, 45(2):178–229, April 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.178;   http://rmp.aps.org/abstract/RMP/
v45/i2/p178_1.

Trimble:1973:SNP

[TR73b] Virginia Trimble and Frederick Reines. The solar neutrino problem — a progress(?) report. Reviews of Modern Physics, 45(1):1–5, January 1973. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.45.1;   http://
rmp.aps.org/abstract/RMP/v45/i1/p1_1.

Trimble:1975:OAC

[Tri75] Virginia Trimble. The origin and abundances of the chemical elements. Reviews of Modern Physics, 47(4):877–976, October 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.877;   http://rmp.aps.org/abstract/RMP/
v47/i4/p877_1.

Tsai:1974:PPB

[Tsa74] Yung-Su Tsai. Pair production and Bremsstrahlung of charged leptons. Reviews of Modern Physics, 46(4):815–851, October 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.46.815;   http://rmp.aps.org/abstract/

RMP/v46/i4/p815_1. See erratum [Tsa77].

Tsai:1977:EPP

[Tsa77] Yung-Su Tsai. Erratum: Pair production and Bremsstrahlung of charged leptons. Reviews of Modern Physics, 49(2):421– 423, April 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.49.421;   http://
rmp.aps.org/abstract/RMP/v49/i2/p421_1. See [Tsa74].

Vajda:1977:AER

[Vaj77] Peter Vajda. Anisotropy of electron radiation damage in metal crystals. Reviews of Modern Physics, 49(3):481–521, July 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.481;   http://rmp.aps.org/abstract/RMP/
v49/i3/p481_1.

VanVleck:1978:QMK

[Van78] J. H. Van Vleck. Quantum mechanics — the key to understanding magnetism. Reviews of Modern Physics, 50(2): 181–189, April 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.50.181;   http://
rmp.aps.org/abstract/RMP/v50/i2/p181_1.

Varma:1976:MVC

[Var76] C. M. Varma. Mixed-valence compounds. Reviews of Modern Physics, 48(2):219–238, April 1976. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.48.219;   http:
//rmp.aps.org/abstract/RMP/v48/i2/p219_1.

Vook:1975:RAP

[VBB^{+}75] F. L. Vook, H. K. Birnbaum, T. H. Blewitt, W. L. Brown, J. W. Corbett, J. H. Crawford, Jr., A. N. Goland, G. L. Kulcinski, M. T. Robinson, D. N. Seidman, F. W. Young, Jr., J. Bardeen, R. W. Balluffi, J. S. Koehler, and G. H. Vineyard (APS Study Group Participants and APS Council Review Committee). Report to the American Physical Society by the study group on physics problems relating to energy technologies: Radiation effects on materials. Reviews of Modern Physics, 47(S3):S1–S44, January 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.47.S1;   http://rmp.aps.org/abstract/
RMP/v47/iS3/pS1_3.

VanVleck:1977:AEL

[VH77] J. H. Van Vleck and D. L. Huber. Absorption, emission, and linebreadths: a semihistorical perspective. Reviews of Modern Physics, 49(4):939–959, October 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.49.939;   http:
//rmp.aps.org/abstract/RMP/v49/i4/p939_1.

Venkataraman:1970:EVC

[VS70] G. Venkataraman and V. C. Sahni. External vibrations in complex crystals. Reviews of Modern Physics, 42(4): 409–470, October 1970. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.42.409;   http:
//rmp.aps.org/abstract/RMP/v42/i4/p409_1. See erratum

[VS71].

Venkataraman:1971:EEV

[VS71] G. Venkataraman and V. C. Sahni. Erratum: External Vibrations in Complex Crystals. Reviews of Modern Physics, 43(4):679, October 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.43.679;   http://
rmp.aps.org/abstract/RMP/v43/i4/p679_1. See [VS70].

Watson:1976:IMR

[Wat76] William D. Watson. Interstellar molecule reactions. Reviews of Modern Physics, 48(4):513–552, October 1976. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.48.513;   http://rmp.aps.org/abstract/RMP/
v48/i4/p513_1.

Wehrl:1978:GPE

[Weh78] Alfred Wehrl. General properties of entropy. Reviews of Modern Physics, 50(2):221–260, April 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.221;   http:
//rmp.aps.org/abstract/RMP/v50/i2/p221_1.

Weinberg:1974:RPG

[Wei74] Steven Weinberg. Recent progress in gauge theories of the weak, electromagnetic and strong interactions. Reviews of Modern Physics, 46(2):255–277, April 1974. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.46.255;   http:
//rmp.aps.org/abstract/RMP/v46/i2/p255_1.

Wheatley:1975:EPS

[Whe75] John C. Wheatley. Experimental properties of superfluid ^{3}He. Reviews of Modern Physics, 47(2):415–470, April 1975. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.47.415;   http://rmp.aps.org/abstract/RMP/
v47/i2/p415_1.

Wilson:1975:RGC

[Wil75] Kenneth G. Wilson. The renormalization group: Critical phenomena and the Kondo problem. Reviews of Modern Physics, 47(4):773–840, October 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.773;   http://
rmp.aps.org/abstract/RMP/v47/i4/p773_1.

Wilson:1977:PLM

[Wil77] Richard Wilson. Physics of liquid metal fast breeder reactor safety. Reviews of Modern Physics, 49(4):893–924, October 1977. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.49.893;   http://rmp.aps.org/abstract/RMP/
v49/i4/p893_1.

Wilson:1979:FFF

[Wil79a] R. R. Wilson. Fantasies of future Fermilab facilities. Reviews of Modern Physics, 51(2):259–273, April 1979. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.259;   http://rmp.aps.org/abstract/RMP/
v51/i2/p259_1.

Wilson:1979:CMB

[Wil79b] R. W. Wilson. The cosmic microwave background radiation. Reviews of Modern Physics, 51(3):433–445, July 1979. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.433;   http://rmp.aps.org/abstract/RMP/
v51/i3/p433_1.

Yennie:1975:HSP

[Yen75] Donald R. Yennie. The hadronic structure of the photon with emphasis on its two-pion constituent. Reviews of Modern Physics, 47(2):311–330, April 1975. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.47.311;   http://
rmp.aps.org/abstract/RMP/v47/i2/p311_1.

Yndurain:1972:RCB

[Ynd72] F. J. Ynduráin. Rigorous constraints, bounds, and relations for scattering amplitudes. Reviews of Modern Physics, 44(4): 645–667, October 1972. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.44.645;   http://
rmp.aps.org/abstract/RMP/v44/i4/p645_1.